### v1

# Wetting Transition on Trees I: Percolation With Clustering

Abstract.
A new “Percolation with Clustering” (PWC) model is introduced, where (the probabilities of) site percolation configurations on the leaf set of a binary tree are rewarded exponentially according to a generic function, which measures the degree of clustering in the configuration. Conditions on such “clustering function” are given for the existence of a limiting free energy and a wetting transition, namely the existence of a non-trivial percolation parameter threshold above and only above which the set of “dry” (open) sites have an asymptotic density. Several examples of clustering functions are given and studied using the general theory. The results here will be used in a sequel paper to study the wetting transition for the discrete Gaussian free field on the tree subject to a hard wall constraint.
Dedicated to the memory of our dear friend and mentor Dima Ioffe. The void left by his sudden passing will remain forever.
1. Introduction
The goal of this manuscript is to study the wetting transition phenomenon for a one-dimensional site percolation-type measure, in which configurations are re-weighted according to some measure of how clustered they are. We shall refer to such measure as percolation with clustering (PWC)
1.1. General theory
For a formal definition, let $\mathbb{T}^{(n)}:=\bigcup_{i=0}^{n}\{0,1\}^{i}$ be the binary tree of depth $n$ rooted at $\mathbf{0}^{(n)}$. While we keep the genealogical relation as usual, for notational convenience it will be useful to the define the age (occasionally level or generation) $|v|$ of vertex (occasionally site) $v\in\mathbb{T}^{(n)}$, as its distance from the leaves, not the root as is often the case. As such, the root (at age $n$) is the oldest vertex, while the leaves (at age 00) are the youngest, and the children of a vertex are one generation younger than their parent. We let $\mathbb{L}_{i}^{(n)}$ denote the set of vertices whose age is $i\in[0,n]$, and for $i\geq|v|$, write $v|_{i}$ for the ancestor of $v$ in $\mathbb{L}_{i}^{(n)}$, The super-index $(n)$ will be occasionally omitted if no confusion is created this way.
Given $n$ and a non-empty subset $A\subseteq\mathbb{L}_{0}^{(n)}$ of the leaves, we let $\mathcal{T}(A)$ be the sub-tree of $\mathbb{T}^{(n)}$ which is generated by $A$. That is, $\mathcal{T}\left(A\right)$ is the subtree of $\mathbb{T}^{(n)}$ which includes all vertices $\{v|_{\ell}:\>v\in A,\ell=0,\dots,n\}$ and all interconnecting edges which are inherited from $\mathbb{T}^{(n)}$. We shall say that two sets $A,A^{\prime}\subset\mathbb{L}_{0}^{(n)}$ are equivalent, and write $A\sim A^{\prime}$, if the corresponding trees $\mathcal{T}(A)$ and $\mathcal{T}(A^{\prime})$ are graph isomorphic. This is clearly an equivalence relation.
The PWC measure is a site-percolation-type measure on the set of leaves $\mathbb{L}_{0}^{(n)}$. As such, it is a distribution over the set of all assignments of “open” and “closed” to each leaf in $\mathbb{L}_{0}^{(n)}$, or equivalently a distribution over subsets $A$ (of open sites) thereof. To adhere to the terminology of the context from which this model arose (see Example 1.2.2 below), we shall refer to open and closed sites as dry and wet respectively and to the set of dry (open) sites as the dry-set. The main ingredient in the definition of this measure is a clustering function, from which the only requirement is
Definition 1.1.
A function $\Phi:2^{\mathbb{L}_{0}^{(n)}}\to\mathbb{R}$ is called clustering if $\Phi(\emptyset)=0$ and $\Phi(A)=\Phi(B)$ whenever $A\sim B$.
With the above at hand, we can thus define:
Definition 1.2.
Let $n\geq 1$, $\Phi:2^{\mathbb{L}_{0}^{(n)}}\to\mathbb{R}$ clustering, and $J\in\mathbb{R}$. The PWC-measure with clustering function $\Phi$ and pinning force $J$ is a probability measure on $2^{\mathbb{L}_{0}^{(n)}}$ which is determined by
$\mathsf{P}_{n}^{\Phi,J}(A):=\frac{1}{\mathbb{Z}_{n}^{\Phi,J}}\,{\rm e}^{J\left|A\right|-\Phi(A)}\quad;\qquad A\subseteq\mathbb{L}_{0}^{(n)}\,,$ (1.1)
where
$\mathbb{Z}_{n}^{\Phi,J}:=\sum_{A\subseteq\mathbb{L}_{0}^{(n)}}{\rm e}^{J\left|A\right|-\Phi(A)}$ (1.2)
is the associated partition function.
We also denote by ${\rm E}_{n}^{\Phi,J}$ the associated expectation. With a bit of abuse, we shall also denote by $A$ the canonical random variable on $2^{\mathbb{L}_{0}^{(n)}}$, so that, e.g. ${\rm P}_{n}^{\Phi,J}(u\in A)$ or ${\rm E}_{n}^{\Phi,J}|A|$ are well defined.
When $\Phi\equiv 0$, one recovers the usual Bernoulli site percolation on $\mathbb{L}_{0}^{(n)}$ with sites being independently open (dry) with probability $p(J)=(1+{\rm e}^{-J})^{-1}$. One should think of $\Phi(A)$ as a function which measures, somehow, the degree of clustering in the set $A$, so that lower value of $\Phi(A)$ means higher degree of clustering in $A$. Then the law $\mathsf{P}_{n}^{\Phi,J}$ can be seen as a modification of the usual percolation distribution in which configurations with a larger degree of clustering are rewarded exponentially.
Note that, contrary to the intuition mentioned earlier, the definition of a clustering function does not inherently relate $\Phi(A)$ to any specific measure of clustering within $A$. While this provides more generality, some results will require such a relationship, which will be achieved by imposing additional constraints on $\Phi$.
The underlying phenomenological question in this work is whether, given a pinning force and a clustering function, the dry set is typically “microscopically” or “macroscopically” large under the PWC measures ${\rm P}_{n}^{\Phi,J}$. Addressing this question necessitates, of course, a formal definition of the above notions of scale. Taking the Physics angle, we fix $J\in\mathbb{R}$ and consider, naturally, not one, but a sequence of clustering function $\{\Phi_{n}\}$ indexed by the height of the tree. We then define the free energy for the corresponding sequence of PWC measures as the limit
$\zeta(J):=\lim_{n\to\infty}\zeta^{\Phi_{n}}_{n}(J)$ (1.3)
where
$\zeta^{\Phi}_{n}(J):=\frac{1}{2^{n}}\ln\mathbb{Z}_{n}^{\Phi,J}\,,$ (1.4)
with $\mathbb{Z}_{n}^{\Phi,J}$ being the partition function from Definition (1.2). While the limit need not exist in general, sufficient conditions under which it does are given below.
Assuming the existence of the limit, we can make a formal sense of the scale of the dry set via the following definition, which is standard in the theory of random surfaces. Again, we adhere to the terminology in the latter context.
Definition 1.3.
Let $J\in\mathbb{R}$ and for each $n$ let $\Phi_{n}$ be a clustering function on $\mathbb{T}_{n}$. The sequence of PWC-measures $\{{\rm P}_{n}^{J,\Phi_{n}}\}$ exhibits complete wetting (a macroscopic dry set) if $\zeta(J)=0$ and otherwise partial wetting (a microscopic dry set).
The standard motivation for the above definition is as follows. Recall that by definition $\Phi_{n}(\emptyset)=0$, so that $\zeta^{\Phi_{n}}_{n}(J)$ is non-negative. If we assume further that $|\Phi_{n}(A)|$ grows at most linearly in $|A|$ uniformly in $n$ (which is natural and will be a consequence of our assumptions in the examples below), then the contribution to the partition function from dry sets whose size is sublinear (in $2^{n}$) is also sublinear on an exponential scale. Consequently, if $\zeta^{\Phi_{n}}_{n}(J)$ is asymptotically positive it must be dominated by dry sets having positive density, i.e. a macroscopic one.
As $\zeta(J)$ is evidently non-decreasing in $J$, it makes sense to state,
Definition 1.4.
The critical pinning force $J^{*}$ associated with the sequence of clustering functions $\{\Phi_{n}\}$ is
$J^{*}=\inf\left\{J\leavevmode\nobreak\ :\leavevmode\nobreak\ \zeta(J)>0\right\}.$ (1.5)
We shall say that $\{\Phi_{n}\}$ exhibits a wetting transition if $J^{*}\in(-\infty,\infty)$.
Next, we wish to impose restrictions on $\{\Phi_{n}\}$ under which the above definitions are proper, namely the free energy exists. These will come in the form of monotonicity with respect to some clustering order on configurations. We shall then derive a necessary and sufficient condition for the existence of a wetting transition. Accordingly, with $u\wedge v$ denoting, customarily, the youngest common ancestor of $u,v\in\mathbb{T}_{n}$, we have,
Definition 1.5.
Given two subsets $A,B\subseteq\mathbb{L}_{0}^{(n)}$ of the same cardinality $\left|A\right|=\left|B\right|$, we say that $A$ is more clustered than $B$ and write $A\prec B$, if there exists a bijection $\sigma:A\to B$, such that
$|u\wedge v|\leq|\sigma_{u}\wedge\sigma_{v}|\quad\forall\,u,v\in A.$ (1.6)
It is easy to see that above relation is a partial order on $2^{\mathbb{L}_{0}{(n)}}$. We can then also have,
Definition 1.6.
A clustering function $\Phi$ is called monotone if
(a) For any $A$ and $B$ of equal cardinality
$A\prec B\ \Rightarrow\ \Phi(A)\leq\Phi(B).$ (1.7)
(b) For any disjoint subsets $A,B\subset\mathbb{L}_{0}$ ,
$\Phi(A\cup B)\leq\Phi(A)+\Phi(B).$ (1.8)
Now, for a clustering function $\Phi$ on $\mathbb{L}_{0}^{(n)}$ and $a_{0}\in[0,2^{n}]$, we define the canonical partition function associated with $\Phi$ as
$\mathbb{W}_{n}(a_{0})\equiv\mathbb{W}^{\Phi}_{n}(a_{0}):=\sum_{\left|A\right|=a_{0}}{\rm e}^{-\Phi(A)}\,,$ (1.9)
so that evidently
$\mathbb{Z}^{\Phi,J}_{n}=\sum_{a_{0}=0}^{2^{n}}{\rm e}^{Ja_{0}}\mathbb{W}^{\Phi}_{n}(a_{0})\,,$ (1.10)
Then given a sequence $\{\Phi_{n}\}$ and $\epsilon\in[0,1]$ dyadic, namely a real number which is equal to $k/2^{m}$ for some integers $k$ and $m$, we also define the canonical free energy as
$\omega(\epsilon):=\lim_{n\to\infty}\omega_{n}^{\Phi_{n}}(\epsilon)\,.$ (1.11)
where
$\omega^{\Phi}_{n}(\epsilon):=\frac{1}{2^{n}}\ln\mathbb{W}_{n}^{\Phi}(\epsilon 2^{n})\,,$ (1.12)
whenever the limit exist.
The next proposition gives a sufficient condition for the existence of the above limit for all dyadics and, consequently, the existence of $\omega$ on all of $[0,1]$. To this end, let us denote by $\mathbb{T}^{(n)}_{{\mathsf{r}}}$ and $\mathbb{T}^{(n)}_{{\mathsf{l}}}$ the two connected components of $\mathbb{T}^{(n)}\setminus{\bf 0}^{(n)}$. These are clearly isomorphic to $\mathbb{T}^{(n-1)}$ and so whenever $A$ is a subset of $\mathbb{L}^{(n)}_{0,{\mathsf{r}}}$ or $\mathbb{L}^{(n)}_{0,{\mathsf{l}}}$, we let $\Phi_{n-1}(A)$ be the value of $\Phi_{n-1}$ on the image of $A$ under such isomorphism.
Proposition 1.7.
Suppose that for all $n\geq 1$, the clustering function $\Phi_{n}$ is monotone and non-negative and that there exists a non-negative sequence $\left\{\gamma_{n}\right\}$, satisfying
$\sum_{n}\gamma_{n}2^{-n}<\infty\,,$ (1.13)
such that for all $n\geq 1$ and all $A\subset\mathbb{L}_{0,{\mathsf{r}}/{\mathsf{l}}}^{(n)}$,
$\Phi_{n}(A)\leq\Phi_{n-1}(A)+\gamma_{n}\,.$ (1.14)
Then the canonical free energy $\omega$ is well defined, finite and concave on all dyadic $\epsilon\in[0,1]$ and therefore can be uniquely continuously extended to all of $[0,1]$.
Henceforth, we will identify the canonical free energy $\omega$ with its continuous extension via the above proposition, whenever its conditions are satisfied. We can now address the existence of the (non-canonical) free energy.
Theorem 1.8.
Let $\{\Phi_{n}\}$ be as in Proposition 1.7. Then the free energy $\zeta$ is well defined, finite and convex on (,)($\infty$,$\infty$)( - ∞ , ∞ ). Furthermore, it is related to the canonical free energy $\omega$ as follows:
$\zeta(J)=\max_{\epsilon\in[0,1]}\left\{J\epsilon+\omega(\epsilon)\right\}.$ (1.15)
A consequence of the above representation is the following characterization of the critical pinning force for the sequence $\{\Phi_{n}\}$.
Corollary 1.9.
Let $\{\Phi_{n}\}$ be as in Proposition 1.7. Then the associated critical pinning force is given by
$J^{*}=-\frac{{\rm d}^{+}}{{\rm d}\epsilon}\omega(0)=-\lim_{\epsilon\downarrow 0}\frac{\omega(\epsilon)}{\epsilon}=-\lim_{\begin{subarray}{c}\epsilon\end{subarray}\downarrow 0}\lim_{n\to\infty}\frac{1}{\epsilon 2^{n}}\ln\mathbb{W}_{n}(\epsilon 2^{n})\,.$ (1.16)
In particular, there is a wetting transition if and only if the right derivative of $\omega$ at 00 is finite.
We note that another conventional way to characterize the scale of the dry set is via the more probabilistic quantity (below $u\in\mathbb{L}_{0}^{(n)}$ is any leaf),
$\rho_{n}(J)=\rho^{\Phi}_{n}(J):=\frac{1}{2^{n}}{\rm E}_{n}^{\Phi,J}|A|={\rm P}_{n}^{\Phi,J}(u\in A)\in[0,1]\,,$ (1.17)
which measures the density of the dry-set. The corresponding asymptotic quantity for a sequence of clustering functions $\{\Phi_{n}\}$ is given by
$\rho(J):=\lim_{n\to\infty}\rho_{n}^{\Phi_{n}}(J)\,,$ (1.18)
assuming again that the limit exists (otherwise, one may take the limit inferior). Complete wetting can then be defined as the regime when $\rho(J)=0$, with partial wetting occurring when the latter is positive.
The link between the two definitions of the wetting regimes is via the relation,
$\zeta_{n}^{\Phi}(J)=\int_{-\infty}^{J}\rho_{n}^{\Phi}(J^{\prime}){\rm d}J^{\prime}\,,$ (1.19)
which can be easily verified by differentiating (1.4). Nevertheless, it is not clear that one can pass to the limit in the above relation. This is resolved by
Corollary 1.10.
Let $\{\Phi_{n}\}$ be as in Proposition 1.7. Then for all $J$ such that the maximum in (1.15) is attained uniquely, the limit in (1.18) exists and
$\rho(J)=\frac{{\rm d}}{{\rm d}J}\zeta(J)=\operatorname*{arg\,max}_{\epsilon\in[0,1]}\left\{J\epsilon+\omega(\epsilon)\right\}.$ (1.20)
Moreover uniqueness in (1.15) holds except for at most countably many values of $J$.
1.2. Examples
Let us now discuss several examples of PWC measures. The first two examples can be thought of as the motivation for the model (see the discussion section following the examples).
1.2.1. Pinned random surfaces
For $n\geq 1$, let $f:\mathbb{R}^{\mathbb{L}^{(n)}_{0}}\to\mathbb{R}_{+}$ be a density function which is invariant under (the restriction to $\mathbb{L}_{0}^{(n)}$ of all isomorphisms of $\mathbb{T}_{n}$. Denote by ${\rm P}^{f}_{n}$ be the probability measure on $\mathbb{R}^{\mathbb{L}_{0}^{(n)}}$ whose density w.r.t. the Lebesgue measure is $f$. One should think of ${\rm P}^{f}_{n}$ as being the law of a random surface $h$ indexed by the leaves, so that $h(u)$ is the height at $u\in\mathbb{L}_{0}^{(n)}$.
For $J\in\mathbb{R}$ we introduce delta pinning at force $J$ by considering the modified law
${\rm P}_{n}^{f,J}({\rm d}h):=\frac{1}{{\rm Z}_{n}^{f,J}}f(h)\prod_{x\in\mathbb{L}_{0}^{(n)}}\big{(}{\rm d}h_{x}+{\rm e}^{J}\delta_{0}({\rm d}h_{x})\big{)}\,,$ (1.21)
with ${\rm Z}_{n}^{f,J}$ chosen as to make the above a probability measure. Interface realizations are thus rewarded exponentially according to number of sites at which the interface is pinned to zero.
Given $A\subset\mathbb{L}_{0}^{(n)}$, let us denote by ${\rm P}_{n}^{f,A}$ the law of $h$ conditional on $\{h|_{A}\equiv 0\}$, so that
${\rm P}^{f,A}_{n}({\rm d}h)=\frac{f(h)}{Z_{n}^{f,A}}\prod_{x\in A^{\rm c}}{\rm d}h_{x}\prod_{x\in A}\delta_{0}({\rm d}h_{x})\;,\quad{\rm Z}_{n}^{f,A}:=\int f(h)\prod_{x\in A^{\rm c}}{\rm d}h_{x}\prod_{x\in A}\delta_{0}({\rm d}h_{x})\,.$ (1.22)
Then, expanding the expression on the right hand side of (1.21), one may rewrite the pinned law ${\rm P}_{n}^{f,J}$ as
${\rm P}_{n}^{f,J}({\rm d}h):=\sum_{A\subset\mathbb{L}_{0}^{(n)}}{\rm e}^{J|A|}\frac{Z_{n}^{f,A}}{{\rm Z}_{n}^{f,J}}\,{\rm P}_{n}^{f,A}({\rm d}h)\,.$ (1.23)
This shows that the pinned law is a convex combination of the conditional measures ${\rm P}_{n}^{f,A}$ with weights given by ${\rm e}^{J|A|}{\rm Z}_{n}^{f,A}/{\rm Z}_{n}^{f,J}$. In particular,
${\rm P}_{n}^{f,J}\big{(}\{0\}^{A}\times(\mathbb{R}\setminus\{0\})^{A^{\rm c}}\big{)}=\frac{1}{{\rm Z}_{n}^{f,J}}{\rm e}^{J|A|}{\rm Z}_{n}^{f,A}\,,$ (1.24)
so that the configuration of pinned sites under ${\rm P}_{n}^{J}$ is a PWC measure with clustering function
$\Phi(A)=\Phi^{f}_{n}(A):=-\ln({\rm Z}_{n}^{f,A}/{\rm Z}_{n}^{f,\emptyset})\quad;\qquad A\subset\mathbb{L}_{0}^{(n)}\,,$ (1.25)
and pinning force $J$.
A particular case where such field $h$ arises, is when one considers the restriction to $\mathbb{L}_{n}^{(0)}$ of a Gibbs gradient field $h^{\prime}$ on $\mathbb{T}_{n}$, whose law is given by
${\rm P}^{V}_{n}({\rm d}h)=\frac{1}{{\rm Z}^{V}_{n}}\exp\Big{\{}-\sum_{x\sim y}V\big{(}h_{x}-h_{y}\big{)}\Big{\}}\prod_{x\in\mathbb{T}_{n}\setminus\{0\}}{\rm d}h_{x}\,\delta_{0}({\rm d}h_{\bf 0})\,,$ (1.26)
where the sum is over all neighboring vertices in $\mathbb{T}_{n}$ and $V:\mathbb{R}\to\mathbb{R}_{+}$ is a smooth, symmetric and convex function. When $V(s)=s^{2}$, the field $h^{\prime}$ is the discrete Gaussian free field (DGFF) on $\mathbb{T}_{n}$. For this particular field, it is known the law of pinned sites under the corresponding measure ${\rm P}^{J,f}_{n}$ is strong FKG [1, Lemma 2.1]. This shows that the corresponding clustering function, we shall denote in this case by
$\Phi^{V}_{n}\equiv\Phi^{f}_{n}\,,$ (1.27)
satisfies the second condition in Definition 1.6 of monotonicity. However the validity of the first condition is not clear.
Nevertheless, for a general $V$, it can be shown by domination over Bernoulli site percolation (cf. [1, Theorem 2.4]) that $\rho^{\Phi}_{n}(J)$ is uniformly positive for all $J>-\infty$. Since the free energy $\zeta$ exists by the strong FKG property of ${\rm P}^{V}_{n}$ (c.f., [4, Lemma 7.8]), we may use (1.19) to conclude that there is no wetting transition in the sense of Definition 1.4, for any $V$ as above.
1.2.2. Pinned random surface above a hard wall
Continuing from the previous example, an interaction of the field $h$ with a, so-called, hard wall placed above $\mathbb{L}_{0}^{(n)}$ at height zero can be modeled by considering the conditional measure
${\rm P}_{n}^{f,+}({\rm d}h):={\rm P}_{n}^{f,+}({\rm d}h\,|\,\Omega^{+})=\frac{{\rm P}_{n}^{f,+}({\rm d}h)1_{\Omega^{+}}(h)}{{\rm P}_{n}^{f,+}(\Omega^{+})}\quad,\qquad\Omega^{+}_{n}:=\mathbb{R}_{+}^{\mathbb{L}_{0}^{(n)}}$ (1.28)
Here, a common consequence of entropy is that the field does not just stay above the wall, but is rather typically pushed away from it. Thus, conditioning introduces a repulsion force. An opposing attraction force can be added, by introducing delta pinning, as before. This leads to considering the law
${\rm P}_{n}^{f,+,J}({\rm d}h):=\frac{1}{{\rm Z}_{n}^{f,+,J}}f(h)\prod_{x\in\mathbb{L}_{0}^{(n)}}\big{(}{\rm d}h_{x}+{\rm e}^{J}\delta_{0}({\rm d}h_{x})\big{)}\quad;\qquad h\in\Omega^{+}_{n}$ (1.29)
Under certain choices of $f$ the forces are in balance, and a non-trivial “competition” between them takes place.
A common interpretation of this model is that of $h$ describing an interface between the liquid and gas phases of some substance put on top of some hard wall. Pinned sites are thus sites at which the gas phase reaches all the way to the wall, making that particular locations dry, hence the terminology of dry and wet sites. Delta pinning then models a force which pulls the gas phase towards the wall. If the gas reaches the wall only at a microscopic region, then the wall is completely wet. Otherwise, wetting is partial. This physical system is the source of the terminology used in this work.
As before, by expanding (1.28) we can express the law of dry sites as
${\rm P}_{n}^{f,+,J}\big{(}\{0\}^{A}\times(\mathbb{R}\setminus\{0\})^{A^{\rm c}}\big{)}=\frac{1}{{\rm Z}_{n}^{f,+,J}}{\rm e}^{J|A|}{\rm Z}_{n}^{f,+,A}$ (1.30)
where
${\rm Z}_{n}^{f,+,A}:=\int_{\Omega^{+}}f(h)\prod_{x\in A^{\rm c}}{\rm d}h_{x}\prod_{x\in A}\delta_{0}({\rm d}h_{x})\,,$ (1.31)
This law is again a PWC measure, this time with clustering function
$\Phi(A)=\Phi^{f,+}_{n}(A):=-\ln({\rm Z}_{n}^{f,+,A}/{\rm Z}_{n}^{f,+,\emptyset})\quad;\qquad A\subset\mathbb{L}_{0}^{(n)}\,,$ (1.32)
and pinning force $J$.
When $h$ is the restriction to $\mathbb{L}_{0}^{(n)}$ of a Gibbs-gradient field $h^{\prime}$ on $\mathbb{T}_{n}$, of the form considered in the previous example, then the corresponding law ${\rm P}_{n}^{f,+}$ is strong FKG (c.f., [5, Lemma 3.1]). Since having a pinned site is a decreasing event, the corresponding PWC-measure (1.30) is also FKG. In particular, this implies that the corresponding clustering function, which we shall denote in this case by
$\Phi^{V,+}_{n}\equiv\Phi^{f,+}_{n}\,,$ (1.33)
satisfies the second condition in Definition 1.6 of monotonicity. Again, whether the first condition thereof holds is not clear.
Nevertheless because of FKG, both the free energy $\zeta$ and the limiting density of pinned sites $\rho$ exist (c.f. [8, Section 6] and [4, Lemma 7.8]). It can be shown, by adapting the proofs when the underlying graph is the two-dimension lattice, that whenever $V$ is Gaussian or Lipschitz that there is a wetting transition in the sense of Definition 1.4 (see, [8, Section 6]).
1.2.3. The zero clustering function
For a sanity check, let us quickly address the $\Phi\equiv 0$ case. As mentioned before, in this case ${\rm P}_{n}^{\Phi,J}$ is the standard Bernoulli site percolation with probability $p(J)=(1+{\rm e}^{-J})^{-1}$ for a site to be dry. In this case, the canonical free energy $\omega(\epsilon)$ is trivially given by the Binomial entropy function
$\mathrm{ent}\left(\epsilon\right):=\epsilon\ln\left(\frac{1}{\epsilon}\right)+(1-\epsilon)\ln\left(\frac{1}{1-\epsilon}\right)$ (1.34)
As the right derivative of $\mathrm{ent}\left(\epsilon\right)$ at 00 is $\infty$, there is no wetting transition. This is in agreement with the (asymptotic) density of dry sites being $p(J)>0$ for all $J>-\infty$.
1.2.4. Clustering via first order branching patterns
Next we turn to examples where clustering is determined directly in terms of the tree $\mathcal{T}(A)$ generated by the dry set $A\subset\mathbb{L}_{0}^{(n)}$. To this end, we define the set of branching points corresponding to $A$ as
$\mathcal{B}(A)=\left\{u\wedge v\leavevmode\nobreak\ \middle|\leavevmode\nobreak\ u,v\in A\right\}\,.$ (1.35)
Observe that $A\subset\mathcal{B}(A)$. The restriction to branching points at level $k\in[0,n]$ is given by
$\mathcal{B}_{k}(A)=\mathcal{B}(A)\cap\mathbb{L}_{k}\,,$ (1.36)
with $b_{k}:=|\mathcal{B}_{k}|$ denoting their number. The first order branching pattern corresponding to $A\subset\mathbb{L}_{0}^{(n)}$ is the sequence
$\underline{b}(A):=(b_{k})_{k=0}^{n}\,.$ (1.37)
We can now state,
Definition 1.11.
Let $n\geq 1$ and $h$, $h_{k}$ for $k\in[0,n]$ be real numbers. Then the function $\Phi\equiv\Phi^{\{h_{k}\},h}_{n}:2^{\mathbb{L}_{0}^{(n)}}\to\mathbb{R}$, defined by setting $\Phi(\emptyset)=0$ and for each $A\subset\mathbb{L}_{0}^{(n)}\setminus\{\emptyset\}$ via
$\Phi(A):=\sum_{k=0}^{n}h_{k}b_{k}+h\quad;\qquad\underline{b}(A)=(b_{k})_{k=0}^{n}\,,$ (1.38)
is called a first order branching clustering function with parameters $(h_{k})_{k=0}^{n}$ and $h$.
As, clearly, $\underline{b}(A)=\underline{b}(A^{\prime})$ for $A\sim A^{\prime}$, the function $\Phi$ is clustering. The next proposition shows that it is also monotone, provided the sequence $(h_{k})_{k\geq 1}$ is such.
Proposition 1.12.
Suppose that $(h_{k})_{k=0}^{n}$ is non-decreasing and that $h\geq h_{n}$, then the corresponding first order branching clustering function $\Phi$ is monotone.
For the remainder of this sub-section we assume that $(h_{k})_{k=0}^{\infty}$ is a given sequence of real numbers and for each $n\geq 1$, we let $\Phi_{n}\equiv\Phi_{n}^{\{h_{k}\},h_{n}}$ be a first order branching clustering function with parameters $(h_{k})_{k=0}^{n}$ and $h=h_{n}$. Turning to the question of wetting, we have
Proposition 1.13.
Suppose that $(h_{k})_{k=0}^{\infty}$ is non-decreasing. Then, the canonical free energy $\omega$ associated with $\{\Phi_{n}\}$ exists if and only if,
$\sum_{k}2^{-k}h_{k}<\infty.$ (1.39)
Next we give sufficient and necessary conditions on the sequence $\{h_{k}\}$ for the existence of a wetting transition.
Proposition 1.14.
Suppose that $(h_{k})_{k=0}^{\infty}$ is non-decreasing and that (1.39) holds. Then there is a wetting transition at $J^{*}\in\left(2\ln 2+h_{0}-\ln\kappa_{1},\,\infty\right)$ if
$\kappa_{1}:=\sum_{k=1}^{\infty}2^{k}{\rm e}^{-h_{k}}<\infty\,,$ (1.40)
Turning to the necessary condition, henceforth we denote the Laplace Transform of sequence $(v_{n})_{v\geq 0}$ by $\hat{v}$. That is,
$\hat{v}(s)=\sum_{n=0}^{\infty}e^{-ns}v_{n}\ ,\ \ s\geq 0\,,$ (1.41)
whenever the series converges absolutely. We shall also write $v^{+}$ for the sequence $(v_{n}\vee 0)_{n\geq 0}$.
Proposition 1.15.
Assume that $(h_{k})_{k=0}^{\infty}$ is non-decreasing and that (1.39) holds. Then there is no wetting transition if
$\lim_{s\downarrow 0}\big{(}\ln\frac{1}{s}-s\widehat{g^{+}}(s)\big{)}=\infty\,,$ (1.42)
where $(g_{k})_{k=0}^{\infty}$ is defined via,
$g_{k}:=h_{k}-(\ln 2)k\quad,\quad k\geq 0\,.$ (1.43)
An immediate consequence of the above is the next corollary which follows by a standard Tauberian argument. It gives an almost complementary condition to that in Proposition 1.14.
Corollary 1.16.
Assume that $(h_{k})_{k=0}^{\infty}$ is a non-decreasing sequence and that (1.39) holds. Then there is no wetting transition if
$\lim_{k\to\infty}k2^{k}{\rm e}^{-h_{k}}=\infty\,,$ (1.44)
or equivalently, if
$\lim_{k\to\infty}(\ln 2)k+\ln k-h_{k}=\infty\,.$ (1.45)
1.2.5. Clustering via second order branching patterns
Next we generalize the previous example, by considering two levels of branching. To this end, if $u,v\in\mathcal{B}(A)$, we shall say that $v$ is a direct branching ancestor of $u$ if $v$ is the youngest ancestor of $u$ which is also a branching point. In this case $u$ is a direct branching descendant of $v$. For $0\leq\ell<k\leq n$, we then let
$\mathcal{B}_{k,\ell}(A):=\big{\{}(u,v)\in\mathcal{B}(A):\>v\text{ direct branching ancestor of }u,\,|u|=\ell,|v|=k\big{\}}\,,$ (1.46)
and set
$b_{k,\ell}:=\begin{cases}|\mathcal{B}_{k,\ell}(A)|&0\leq\ell<k\leq n\,,\\ {1\hskip-3.0pt\rm{I}}_{\{\max_{\mathcal{B}(A)}|v|\}}(\ell)&0\leq\ell<k=n+1\,.\end{cases}$ (1.47)
Thus, $b_{k,\ell}$ is the number of branching points at level $\ell$ whose direct branching ancestor is in level $k$, where the oldest branching point: $\operatorname*{arg\,max}_{v\in\mathcal{B}(A)}|v|$, is thought of as having a branching ancestor at level $n+1$. The second order branching pattern corresponding to $A\subset\mathbb{L}_{0}^{(n)}$ is the triangular array:
[…]
where we use the same notation from the previous example. Observe that the first order branching pattern elements $b_{k}$ can be easily reconstructed from the second order pattern via,
$b_{i}(A)=\sum_{k=i+1}^{n+1}b_{k,i}=\frac{1}{2}\sum_{\ell=0}^{i-1}b_{i,\ell}\,.$ (1.49)
We can now define the second order branching clustering function.
Definition 1.17.
[…]
$\Phi(A):=\sum_{0\leq\ell<k\leq n+1}h_{k,\ell}b_{k,\ell}(A)+h\;;\quad\underline{b}(A)=(b_{k,\ell})_{0\leq\ell<k\leq n+1}\,,$ (1.50)
is called a second order branching clustering function with parameters $h$ and $(h_{k,\ell})_{0\leq\ell<k\leq n+1}$.
As in the previous example, the clustering property of $\Phi$ follows essentially by definition. Monotonicity, on the other hand, is inherited from that of the triangular array. To this end, we shall say that triangular array $\underline{h}=\left\{h_{k,\ell}\right\}$ is non-decreasing if
$k\leq k^{\prime}\ {\rm and}\ \ell\leq\ell^{\prime}\ \Rightarrow h_{k,\ell}\leq h_{k^{\prime},\ell^{\prime}}\,.$ (1.51)
Proposition 1.18.
Let $n\geq 1$. Suppose that $\{h_{k,\ell}\}_{0\leq\ell<k\leq n+1}$ is a non-decreasing triangular array and that $h\geq h_{n+1,n}$. Then a second order branching clustering function with parameters $\{h_{k,\ell}\}_{0\leq\ell<k\leq n+1}$ and $h$ is a monotone clustering function.
Turning to the question of wetting, as in the previous example we suppose that an infinite triangular array $(h_{k,\ell})_{0\leq\ell<k}$ is given, and for $n\geq 1$ let $\Phi_{n}\equiv\Phi_{n}^{\{h_{k,\ell}\},h_{n+1,n}}$ be a second order branching clustering function with parameters $(h_{k,\ell})_{0\leq\ell<k\leq n+1}$ and $h=h_{n+1,n}$.
Proposition 1.19.
Suppose that $(h_{k,\ell})_{0\leq\ell<k}$ is non-decreasing. Then, the free energy exists if
$\sum_{k}2^{-k}h_{k,k-1}<\infty\,.$ (1.52)
For a sufficient condition on $\{h_{k,\ell}\}$ for the existence of a wetting transition
Proposition 1.20.
Suppose that $(h_{k,\ell})_{0\leq\ell<k}$ is non-decreasing and that (1.52) holds. Then, there is a wetting transition at $J^{*}\in\big{(}2\ln 2-2e^{-1}\kappa_{2},\,\infty\big{)}$ if
$\kappa_{2}:=\sup\limits_{k}\sum\limits_{\ell=0}^{k-1}2^{\ell}e^{-h_{k,\ell}}<\infty.$ (1.53)
Turning to the necessary condition, if $v=(v_{n_{1},n_{2}})_{n_{1},n_{2}\geq 0}$ is a doubly indexed sequence, then we denote its (bi-variate) Laplace Transform by $\hat{v}(s_{1},s_{2})$. That is,
$\hat{v}\left(s_{1},s_{2}\right)=\sum_{n_{1}=0}^{\infty}\sum_{n_{2}=0}^{\infty}e^{-s_{1}n_{1}-s_{2}n_{2}}\,v_{n_{1},n_{2}}\,,$ (1.54)
whenever that the series converges absolutely.
Proposition 1.21.
Suppose that $(h_{k,\ell})_{0\leq\ell<k}$ is non-decreasing and that (1.52) holds. Then, there is no wetting transition if
$\lim_{s\downarrow 0}\big{(}\ln{\frac{1}{s}}-2s^{2}\widehat{g^{+}}\left(s,2s\right)\big{)}=\infty\,,$ (1.55)
where $\{g_{\ell,d}\}$ is a doubly indexed sequence defined via,
$g_{\ell,d}:=\left\{\begin{array}[]{lll}h_{\ell+d,\ell}-(\ln{2})\ell&&l\geq 0,d\geq 1\,,\\ 0&&l\geq 0,d=0\,.\end{array}\right.$ (1.58)
By making stronger assumptions on the structure of $\{h_{k,\ell}\}$, we have the following simpler necessary condition.
Corollary 1.22.
Suppose that $(h_{k,\ell})_{0\leq\ell<k}$ is non-decreasing, that (1.52) holds and that
$h_{\ell+d,\ell}=h^{(1)}_{\ell}+h^{(2)}_{d}\;;\quad\ell,d\geq 0\,,$ (1.59)
for two non-negative sequences $(h^{(1)}_{k})_{k\geq 0}$ and $(h^{(2)}_{k})_{k\geq 0}$. Then there is no wetting transition if
$\lim_{k\to\infty}\,(\ln 2)k+\ln{k}-h_{k}^{\left(1\right)}-h_{\lfloor k/2\rfloor}^{\left(2\right)}=\infty\,.$ (1.60)
Remark 1.23.
When $h_{k,\ell}\equiv h_{\ell}$ does not depend on $k$, the function $\Phi_{n}$ becomes a first order branching clustering functions with $(h_{\ell})_{\ell\geq 0}$ and $h=h_{n}$. In this case we may take $h^{(2)}_{d}\equiv 0$ and $h_{\ell}^{(1)}=h_{\ell}$, in decomposition (1.59). Then it is not difficult to see that conditions (1.55) and (1.60) identify with conditions (1.42) and (1.45) respectively. We also note that, the assumption on the positivity of the two sequences in the decomposition can be weakened to requiring that $(h^{(1)}_{k})^{-},(h^{(2)}_{k})^{-}=o(k\log k)$, with minor changes to the proof.
1.2.6. Clustering via capacity
Lastly we provide another, somehow surprising, example. Let $\mathcal{C}=\left\{\mathsf{C}_{e}\right\}$ be a family of conductances, such that $\mathsf{C}_{e}=\mathsf{C}_{\left|e\right|}$, where $\left|e\right|:=\left|u\right|$ if $e=e(u)$ is the edge adjacent to $u$ in the direction of the root $\mathsf{0}$. Recall that the capacity of a set $A$ is
${\rm CAP}_{\mathcal{C}}(A):=\min_{f(\mathsf{0})=1,\ f_{A}=0}\sum_{u\neq\mathsf{0}}\mathsf{C}_{e(u)}\left(\nabla_{e(u)}f\right)^{2}.$ (1.61)
A motivation for using capacity as a clustering function can be seen by recalling the representation of the former via escape probabilities of a random walk on $\mathbb{T}^{(n)}$. Think of $\mathcal{C}_{e}$ in terms of a jump rate across an un-oriented bond $e$. This gives rise to a continuous time random walk on $\mathbb{T}^{(n)}$. Let $S_{t}$ be its discrete time jump chain, and let $\mathbb{P}^{\rm RW}_{u}$ be the distribution of $S$ on $\mathbb{T}^{(n)}$ starting at $u$. Finally let $\tau_{v}$ be the first hitting/return time to $v$; $\tau_{v}=\min\left\{t>0\leavevmode\nobreak\ :\leavevmode\nobreak\ S_{t}=v\right\}$ and set $\tau_{A}=\min_{v\in A}\tau_{v}$. Then,
${\rm CAP}_{\mathcal{C}}(A)=\mathbb{P}_{\mathsf{0}}^{\rm RW}\left(\tau_{A}<\tau_{\mathsf{0}}\right)\cdot\sum_{z\sim\mathsf{0}}\mathsf{C}_{e(z)}=\mathsf{C}_{0}\sum_{v\in A}\mathbb{P}_{v}^{\rm RW}\left(\tau_{\mathsf{0}}<\tau_{A}\right)$ (1.62)
In particular, if $A$ and $B$ are disjoint, then
$\begin{split}\mathbb{P}_{\mathsf{0}}^{\rm RW}\left(\tau_{A\cup B}<\tau_{\mathsf{0}}\right)&=\mathbb{P}_{\mathsf{0}}^{\rm RW}\left(\tau_{A}<\tau_{\mathsf{0}}\wedge\tau_{B}\right)+\mathbb{P}_{\mathsf{0}}^{\rm RW}\left(\tau_{B}<\tau_{\mathsf{0}}\wedge\tau_{A}\right)\\ &\leq\mathbb{P}_{\mathsf{0}}^{\rm RW}\left(\tau_{A}<\tau_{\mathsf{0}}\right)+\mathbb{P}_{\mathsf{0}}^{\rm RW}\left(\tau_{B}<\tau_{\mathsf{0}}\right),\end{split}$
which implies (1.8).
It takes a bit more work to show,
Proposition 1.24.
Let $\Phi(A):={\rm CAP}_{\mathcal{C}}(A)$. Then $\Phi$ is a monotone clustering function.
Turning to wetting. As it turns out, there is no wetting transition in this case.
Proposition 1.25.
There exists $C_{0}<\infty$ such that for all $n$ and $A\subset\mathbb{L}_{0}^{(n)}$,
${\rm CAP}_{\mathcal{C}}(A)\leq C_{0}\left|A\right|.$ (1.63)
Consequently, there is no wetting transition in this case.
1.3. Discussion and open questions
Let us make some concluding remarks. First, aside from intrinsic interest, the motivation behind studying PWC measures driven by first and second order branching clustering functions is that they can be used to study the wetting transition in the more physical cases of pinned random surfaces. Indeed, in a sequel paper [3] we shall study in more detail the PWC measure of Example 1.2.2 in the case that the field $h$ is the restriction to the leaves of the discrete Gaussian free field $h^{\prime}$ on $\mathbb{T}_{n}$, namely when $V(s)=s^{2}$ in (1.26).
The strong FKG property of the law of $h^{\prime}$ which still holds under the conditioning on positivity at all leaves, will then allow us to show that
$\Phi^{V}_{n}(A)=\Phi^{\{h_{k,\ell},h_{n+1,n}\}}_{n}(A)+O(1)\quad;\qquad A\subset\mathbb{L}_{0}^{(n)}\,,$ (1.64)
where $\Phi^{V}_{n}$ is the clustering function in Example 1.2.2, while $\Phi_{n}^{\{h_{k,\ell}\},h_{n+1,n}}$ is the second order branching clustering function of Example 1.2.5 with
$h_{k,\ell}:=(\ln 2)\ell+\frac{3}{2}\ln^{+}(\ell\wedge(k-\ell))\;;\quad 0\leq\ell\leq k\leq n+1\,.$ (1.65)
As the conditions in Proposition 1.20 hold, the PWC measures associated with $\Phi_{n}^{\{h_{k,\ell}\},h_{n+1,n}}$ exhibit a wetting transition.
By monotonicity of the canonical free energy in $\Phi$, domination (1.64) and Corollary 1.9, this implies that there is a wetting transition also for the PWC measures associated with $\Phi^{V}_{n}$. This gives an alternative proof to the Caputo-Velenik argument for the existence of a wetting transition for the DGFF on the tree subject to a hard wall (cf. [2]), which is more constructive in nature. We note that one can have a similar comparison of $\Phi_{n}^{V}$ with a first order branching cluster functions w.r.t. the sequence $h_{k}=(\ln 2)k+\ln k$. The trouble is that this case falls exactly “between” the necessary and sufficient conditions and Example 1.2.4 cannot be used to rule out or in the existence of a wetting transition.
Many questions concerning the PWC model and Examples 1.2.4 and 1.2.5 in particular remain. Let us now mention some of them. When there is a wetting transition at $J^{*}\in(-\infty,\infty)$, a natural question is whether $\rho$ is continuous at $J^{*}$ or not, i.e. what the order of the wetting transition is. This question is directly related to the curvature of $\omega$ at 00, as can be seen by (1.15). Other questions concern the stochastic geometry of pinned sites in the complete and partial wetting regime. These can be phrased in term of the fractal nature of the dry-set (e.g. typical number of dry points in a ball of radius $\alpha n$ in graph-distance, or correlations between the events that two sites are pinned.
Both answers in the generality of the basic theory and for Examples 1.2.4 and 1.2.5 are desirable. The latter, because relation (1.64) will likely allow us to deduce conclusions about the dry-set in the pinned random surfaces case, where many of the above questions remain open.
Lastly, while we treat the binary tree here, a more interesting underlying graph would naturally be a $d$-dimensional lattice. While these two graphs seem very different, a tree-like structure can be extracted from a $d$-dimensional box, by, e.g., repeatedly partitioning the box to $2^{d}$ sub-boxes, each having half the side-length of the original one. As in the tree case, this may provide additional tools for studying the wetting transition in the case of $d$-dimensional random surfaces, and in particular the $d=2$ case, where the current theory is quite lacking. We believe that the general theory, and Examples 1.2.4 and 1.2.5 could be adapted to the lattice case, as well as the comparison between the latter and the DGFF when $d=2$, but we leave this to a future work.
Paper outline
The remainder of the paper is organized as follows. In Section 2 we provide proofs for the general theory of wetting as stated in the introduction. Sections 3 and 4 include the proofs for the statements concerning PWC measures driven by first and second order branching clustering functions, respectively. Section 5 treats the case of clustering function defined in terms of capacity.
2. General theory of wetting
In this section we provide proofs for the statements of the general theory developed in the Subsection 1.1
Proof of Proposition 1.7.
Fix dyadic $\epsilon_{r},\epsilon_{\ell}\in(0,1)$ and set $\epsilon=\frac{1}{2}\left(\epsilon_{r}+\epsilon_{\ell}\right)$. Clearly both $\epsilon_{r}2^{n-1}$ and $\epsilon_{\ell}2^{n-1}$ are in $\mathbb{N}$ for all $n$ sufficiently large. In the latter case if $A_{r}$ is a subset of $\mathbb{L}_{0,{\mathsf{r}}}^{(n)}$ of cardinality $\epsilon_{r}2^{n-1}$ and $A_{\ell}$ is a subset of $\mathbb{L}_{0,{\mathsf{l}}}^{(n)}$ of cardinality $\epsilon_{\ell}2^{n-1}$, then $A=A_{r}\cup A_{\ell}$ has cardinality $\epsilon 2^{n}$. Therefore, by monotonicity and in view of assumption (1.14),
$\mathbb{W}_{n}(\epsilon)\geq{\rm e}^{-2\gamma_{n}}\mathbb{W}_{n-1}(\epsilon_{r})\mathbb{W}_{n-1}(\epsilon_{\ell}).$ (2.1)
Take $\epsilon=\epsilon_{\ell}=\epsilon_{r}$. Then standard sub-additivity arguments imply that $\omega(\epsilon)$ is well defined and satisfies
$\mathrm{ent}\left(\epsilon\right)\geq\omega(\epsilon)\geq\omega_{n}(\epsilon)-2\sum_{n+1}^{\infty}\gamma_{k}2^{-k}>-\infty.$ (2.2)
where $\mathrm{ent}\left(\epsilon\right)$ is as in (1.34). Going back to (2.1) we infer that under assumption (1.14),
$\omega\left(\frac{\epsilon_{{\mathsf{r}}}+\epsilon_{{\mathsf{l}}}}{2}\right)\geq\frac{1}{2}\left(\omega(\epsilon_{{\mathsf{r}}})+\omega(\epsilon_{{\mathsf{l}}})\right),$
for general dyadic $\epsilon_{{\mathsf{r}}/{\mathsf{l}}}\in[0,1]$. By iterations, this is concavity. ∎
Next, we prove
Proof of Theorem 1.8.
Convexity and hence continuity of the right hand side of (1.15) follows by concavity of $\omega$. At the same time, abbreviating $\omega_{n}\equiv\omega_{n}^{\Phi_{n}}$, we have
$\zeta_{n}(J)=\frac{1}{2^{n}}\ln\sum_{\epsilon\in[0,1]}{\rm e}^{2^{n}\left[J\epsilon+\omega_{n}(\epsilon)\right]}$
Bounding the sum using the maximal summand we have
$\max_{\epsilon\in[0,1]}\left\{J\epsilon+\omega_{n}(\epsilon)\right\}\leq\zeta_{n}(J)\leq\max_{\epsilon\in[0,1]}\left\{J\epsilon+\omega_{n}(\epsilon)\right\}+\frac{n\ln 2}{2^{n}},$ (2.3)
with $\epsilon$ in the summation and maximum restricted to $2^{-n}\mathbb{N}_{0}$. Approximating an argmax of (1.15) via a dyadic $\epsilon$ and using the convergence of $\omega_{n}$ to $\omega$ shows that the limit inferior of the lower bound in (2.3) is at least the right hand side (1.15). On the other hand thanks to the second inequality in (2.2) we may replace $\omega_{n}$ by $\omega$ in the upper bound in (2.3) at the cost of an error that tends to 00 with $n$. This shows that the limit superior of the upper bound also tends to the right hand side of (1.15). ∎
Proof of Corollary 1.9.
Since $\omega$ is concave and $\omega(0)=0$, if $(-J)\geq\frac{{\rm d}^{+}}{{\rm d}\epsilon}\omega(0)$ then for all $\epsilon\in[0,1]$
$\omega(\epsilon)\leq(-J)\epsilon\,.$ (2.4)
Also, by definition of the right derivative, if $(-J)<\frac{{\rm d}^{+}}{{\rm d}\epsilon}\omega(0)$ then in a right neighborhood of 00 we must have
$\omega(\epsilon)>(-J)\epsilon$ (2.5)
Proof of Corollary 1.10.
Let $J$ be such that the maximum in (1.15) is attained uniquely at $\epsilon_{0}\in[0,1]$. Then for all $\delta>0$, by continuity, the same maximum restricted further to $(\epsilon_{0}-\delta,\epsilon_{0}+\delta)^{\rm c}$ must be smaller than $\zeta(J)-\eta$ for some $\eta>0$. In particular, bounding the sum using the maximal summand and appealing to (2.2), we have
$\begin{split}\zeta_{n}(J;\delta):&=\frac{1}{2^{n}}\ln\sum_{|2^{-n}|A|-\epsilon_{0}|\geq\delta}{\rm e}^{J\left|A\right|-\Phi(A)}=\frac{1}{2^{n}}\ln\sum_{|\epsilon-\epsilon_{0}|\geq\delta}{\rm e}^{2^{n}\left[J\epsilon+\omega_{n}(\epsilon)\right]}\\ &\leq\max_{|\epsilon-\epsilon_{0}|\geq\delta}\left\{J\epsilon+\omega_{n}(\epsilon)\right\}+\frac{n\ln 2}{2^{n}}\leq\zeta(J)-\eta+2\sum_{n+1}^{\infty}\gamma_{k}2^{-k}+\frac{n\ln 2}{2^{n}}\end{split}$ (2.6)
Since
$\frac{1}{2^{n}}\ln{\rm P}_{n}^{\Phi,J}\Big{(}\big{|}2^{-n}|A|-\epsilon_{0}\big{|}\geq\delta\Big{)}=\zeta_{n}(J;\delta)-\zeta_{n}(J)\,,$ (2.7)
we thus get
$\limsup_{n\to\infty}\,\ln{\rm P}_{n}^{\Phi,J}\Big{(}\big{|}2^{-n}|A|-\epsilon_{0}\big{|}\geq\delta\Big{)}\leq-\eta\,,$ (2.8)
which implies exponential convergence in probability to $\epsilon_{0}$ of $2^{-n}|A|$, and thanks to the boundedness of this quantity also in $L^{1}$.
In the same vein, for any $\Delta>0$,
$\zeta(J)+\epsilon_{0}\Delta\leq\zeta(J+\Delta)\leq\max\Big{\{}\zeta(J)+(\epsilon_{0}+\delta)\Delta\,,\,\zeta(J)-\eta+\Delta\Big{\}}\,,$ (2.9)
Subtracting $\zeta(J)$ and then dividing by $\Delta$ and finally taking $\Delta\downarrow 0$ follows by $\delta\to 0$, shows that the right derivative of $\zeta$ exists at $J$ and equal to $\epsilon_{0}$. A symmetric argument shows the same for the left derivative.
Lastly, since $J\epsilon+\omega(\epsilon)$ is concave, if it has more than one maximizer on $[0,1]$, it must be a constant on some sub-interval of $[0,1]$, which implies that $\omega(\epsilon)$ is linear on that domain with slope $-J$. Since these sub-intervals must be disjoint for different $J$-s, there could be at most countably many of them. ∎
3. Clustering via first order branching patterns
In this section we provide proofs for the various statements in Subsection 1.2.4.
3.1. Monotoncity
Proof of Proposition 1.12.
For $k=0,\dots,n$, let $\beta_{k}=\sum_{0}^{k}b_{j}$ be the number of branching points in all the generations starting from generation $k$ (notice that we refer to the leaves as branching points). Alternatively, $b_{k}=\beta_{k}-\beta_{k-1}$. Of course, if $\left|A\right|=N+1$, then $\beta_{n}(A)=\left|\mathcal{B}(A)\right|=N+N+1=2N+1$. Summing by parts,
$\Phi_{n}(A)=\sum_{0}^{n}h_{k}b_{k}+h=Nh_{n}-\sum_{k=0}^{n-1}(h_{k+1}-h_{k})\beta_{k}(A)+h.$ (3.1)
First we will show (1.7). Let $A\prec C$ and let $\sigma:A\rightarrow C$ be the bijection from the definition of $A\prec C$. We will prove that $\beta_{k}(A)\geq\beta_{k}(C)$ by induction on $|A|=|C|=N+1$. If $N=0$ then $|\mathcal{B}(A)|=|\mathcal{B}(C)|=0$ so $\beta_{k}(A)\geq\beta_{k}(C)$ trivially for all $k=0,\dots,n$. Assume that $\beta_{k}(A)\geq\beta_{k}(C)$ for all $A\prec C$ of cardinality $N+1$. Let $A\prec C$, $|A|=|C|=N+2$. Label $A=\{1,\dots,N+2\}$, $C=\{\sigma_{1},\dots,\sigma_{N+2}\}$. Define $m_{i}=\min_{i\neq j}|i\wedge j|$. Without loss of generality, assume $|1\wedge 2|=m_{1}=\min_{i}m_{i}$. Consider now $\hat{A}=A\setminus\left\{1\right\}$, $\hat{C}=C\setminus\left\{\sigma_{1}\right\}$. By restricting $\sigma$ to $\hat{A}$ we get $\hat{A}\prec\hat{C}$. By the induction hypothesis $\beta_{k}(\hat{A})\geq\beta_{k}(\hat{C})$ for all $k=0,\dots,n$. Since adding one leaf to $\hat{A}$ implies a single additional branching point at generation $m_{i}$ we get
$\beta_{k}(A)=\begin{cases}\beta_{k}(\hat{A})+1&k\geq m_{i}\\ \beta_{k}(\hat{A})&k<m_{i}\end{cases}$
Denote $m^{C}_{1}=\min_{j\neq 1}|\sigma_{1}\wedge\sigma_{j}|$. Then
$\beta_{k}(C)=\begin{cases}\beta_{k}(\hat{C})+1&k\geq m^{C}_{i}\\ \beta_{k}(\hat{C})&k<m^{C}_{i}\end{cases}$
Since $m_{1}\leq m^{C}_{1}$,
$\forall k=0,\dots,n\quad\beta_{k}(A)\geq\beta_{k}(C).$
Since $h_{k}$ is non-decreasing, this implies (1.7). Moving on to (1.8), let $A,C\subseteq\mathbb{L}_{0}$ be disjoint, $A\cap C=\emptyset$. Then
$|A\cup C|=|A|+|C|\Longrightarrow\beta_{n}(A\cup C)=\beta_{n}(A)+\beta_{n}(C)+1.$
First, consider the case where $\mathcal{B}(A)\cap\mathcal{B}(C)=\emptyset$. Because $\mathcal{B}(A),\mathcal{B}(C)\subseteq\mathcal{B}(A\cup C)$ we get $\mathcal{B}(A\cup C)=\mathcal{B}(A)\cup\mathcal{B}(C)\cup\left\{z\right\}$ so:
$b_{|z|}(A\cup C)=b_{|z|}(A)+b_{|z|}(C)+1$
$\forall k\in\left\{0,\dots,n\right\},k\neq|z|\quad b_{k}(A\cup C)=b_{k}(A)+b_{k}(C)$
Therefore
$\forall k=0,\dots,n\quad\beta_{k}(A\cup C)\geq\beta_{k}(A)+\beta_{k}(C).$
Now suppose that $\mathcal{B}(A)\cap\mathcal{B}(C)\neq\emptyset$. Define
$\mathcal{B}_{AC}:=\mathcal{B}(A)\cap\mathcal{B}(C),\ |\mathcal{B}_{AC}|=N>0$
$\mathcal{B}_{A\Delta C}:=\mathcal{B}(A\cup C)\setminus(\mathcal{B}(A)\cup\mathcal{B}(C))$
since $\mathcal{B}(A),\mathcal{B}(C)\subseteq\mathcal{B}(A\cup C)$,
$|\mathcal{B}(A)\cup\mathcal{B}(C)|$ $=|\mathcal{B}(A)|+|\mathcal{B}(C)|-|\mathcal{B}(A)\cap\mathcal{B}(C)|$
$=\beta_{n}(A)+\beta_{n}(C)-N$
and
$|\mathcal{B}_{A\Delta C}|$ $=|\mathcal{B}(A\cup C)|-|\mathcal{B}(A)\cup\mathcal{B}(C)|$
$=\beta_{n}(A\cup C)-(\beta_{n}(A)+\beta_{n}(C)-N)$
$=N+1\,.$
Every $z\in\mathcal{B}_{AC}$ is counted twice in $\beta_{k}(A)+\beta_{k}(C)$, for $k\geq|z|$, but only once in $\beta_{k}(A\cup C)$. On the other hand, $z\in\mathcal{B}_{A\Delta C}$ are only counted in $\beta_{k}(A\cup C)$, $k\geq|z|$. Hence, we get (1.8) by constructing a bijection $\tau:\mathcal{B}_{A\Delta C}\rightarrow\mathcal{B}_{AC}\cup\left\{t\right\}$ satisfying:
$\forall z\in\mathcal{B}_{A\Delta C}\quad|\tau(z)|\geq|z|\text{ or }\tau(z)=t$
where $t$ is a placeholder node.
Label $\mathcal{B}_{A\Delta C}=\left\{1,\dots,N+1\right\}$.
Lemma 3.1.
Any $x\in\mathcal{B}_{AC}$ has no ancestor in $\mathcal{B}_{A\Delta C}$.
Proof.
Assume by way of contradiction that $\exists x\in\mathcal{B}_{AC},z\in\mathcal{B}_{A\Delta C}$ such that $z$ is an ancestor of $x$. By the definitions of $\mathcal{B}_{AC}$ and $\mathcal{B}_{A\Delta C}$ there exist $u_{1},u_{2},u_{3}\in A$, $v_{1},v_{2},v_{3}\in C$ such that $x=u_{1}\wedge u_{2}=v_{1}\wedge v_{2}$ and $z=u_{3}\wedge v_{3}$. We call a path $P=\langle s_{1},\dots,s_{k}\rangle$ in $\mathbb{T}^{(n)}$ an ascending path if $|s_{i+1}|>|s_{i}|$ for all $i=1,\dots,k-1$. Since $z$ is an ancestor of $x$, there exists an ascending path $P_{1}=\langle x,\dots,z\rangle$. Denote by $z_{1},z_{2}$ the direct children of $z$. Without loss of generality, assume $u_{3}$ is a descendant of $z_{1}$, $v_{3}$ is a descendant of $z_{2}$, and that $P_{1}=\langle x,\dots,z_{1},z\rangle$. Let $P_{2}=\langle v_{1},\dots,x\rangle$ be the ascending path from $v_{1}$ to $x$, and $P_{3}=\langle v_{3},\dots,z_{2},z\rangle$ be the ascending path from $v_{3}$ to $z$. Since $z_{1}\neq z_{2}$ and this is a tree, $\left\{P_{2}\cdot P_{1}\right\}\cap\left\{P_{3}\right\}=\left\{z\right\}$, i.e. the simple paths from $v_{1}$ and $v_{3}$ to the root intersect for the first time at $z$ so $z=v_{1}\wedge v_{3}$. This means that $z\in\mathcal{B}(C)$ and therefore $z\notin\mathcal{B}_{A\Delta C}$ which is a contradiction. ∎
Lemma 3.2.
$\forall z_{1}\neq z_{2}\in\mathcal{B}_{A\Delta C},\quad x=z_{1}\wedge z_{2}\in\mathcal{B}_{AC}.$
Proof.
$z_{1}=u_{1}\wedge v_{1}$, $z_{2}=u_{2}\wedge v_{2}$ where $u_{1},u_{2}\in A$ and $v_{1},v_{2}\in C$. Then $x$ is an ancestor of $v_{1}$ and $v_{2}$ and they are in different branches starting at $x$ (symmetrically, also for $u_{1}$ and $u_{2}$). Hence $x\in\mathcal{B}_{AC}$. ∎
Lemma 3.3.
$\forall x_{1}\neq x_{2}\in\mathcal{B}_{AC},\quad x=x_{1}\wedge x_{2}\in\mathcal{B}_{AC}.$
Proof.
By definition of $\mathcal{B}_{AC}$, $x_{1}=u_{1}\wedge u_{2}=v_{1}\wedge v_{2}$ and $x_{2}=u_{1}^{\prime}\wedge u_{2}^{\prime}=v_{1}^{\prime}\wedge v_{2}^{\prime}$ where $u_{1},u_{2},u_{1}^{\prime},u_{2}^{\prime}\in A$ and $v_{1},v_{2},v_{1}^{\prime},v_{2}^{\prime}\in C$. If $x_{1}=x_{2}$ then the lemma is trivial. Otherwise, $x_{1}\neq x_{2}$. Since $x_{1}\neq x_{2}$ and $x$ is the lowest (closest to the leaves) common ancestor of $x_{1}$ and $x_{2}$, $\left\{u_{1},u_{2},v_{1},v_{2}\right\}$ and $\left\{u_{1}^{\prime},u_{2}^{\prime},v_{1}^{\prime},v_{2}^{\prime}\right\}$ are in different branches of $x$ so $x=u_{1}\wedge u_{1}^{\prime}=v_{1}\wedge v_{1}^{\prime}\in\mathcal{B}_{AC}$. ∎
Lemma 3.4.
$\forall x\in\mathcal{B}_{AC},\exists z_{1}\neq z_{2}\in(\mathcal{B}_{A\Delta C}\cup\mathcal{B}_{AC})\setminus\left\{x\right\}:x=z_{1}\wedge z_{2}$
Proof.
Let $x\in\mathcal{B}_{AC}$. $x=u_{1}\wedge u_{2}=v_{1}\wedge v_{2}$ for $u_{1},u_{2}\in A$ and $v_{1},v_{2}\in C$. Denote $x_{1},x_{2}$ the direct children of $x$. Without loss of generality, $x_{1}$ is an ancestor of $u_{1},v_{1}$ and $x_{2}$ is an ancestor of $u_{2},v_{2}$. Then $|u_{1}\wedge v_{1}|\leq|x_{1}|<|x|$, $|u_{2}\wedge v_{2}|\leq|x_{2}|<|x|$. For $i=1,2$, denote $z_{i}^{(0)}=u_{i}\wedge v_{i}$. If $z_{i}^{(0)}\notin(\mathcal{B}_{A\Delta C}\cup\mathcal{B}_{AC})\setminus\left\{x\right\}$ then either $z_{i}^{(0)}\in\mathcal{B}(A)\setminus\mathcal{B}(C)$ or $z_{i}^{(0)}\in\mathcal{B}(C)\setminus\mathcal{B}(A)$. Without loss of generality assume that $z_{i}^{(0)}\in\mathcal{B}(A)\setminus\mathcal{B}(C)$. Then there exists $u_{i}^{(1)}\in A$ such that $z_{i}^{(1)}=u_{i}^{(1)}\wedge v_{i}$ is a descendant of $z_{i}$ satisfying $|z_{i}^{(1)}|<|z_{i}^{(0)}|$. If $z_{i}^{(1)}\notin(\mathcal{B}_{A\Delta C}\cup\mathcal{B}_{AC})\setminus\left\{x\right\}$ then either $z_{i}^{(1)}\in\mathcal{B}(A)\setminus\mathcal{B}(C)$ or $z_{i}^{(1)}\in\mathcal{B}(C)\setminus\mathcal{B}(A)$. Without loss of generality assume that $z_{i}^{(1)}\in\mathcal{B}(A)\setminus\mathcal{B}(C)$. Then there exists $u_{i}^{(2)}\in A$ such that $z_{i}^{(2)}=u_{i}^{(2)}\wedge v_{i}$ is a descendant of $z_{i}^{(1)}$ satisfying $|z_{i}^{(2)}|<|z_{i}^{(1)}|$. We can continue this process, choosing $z_{i}^{(0)},\dots,z_{i}^{(k)}$ such that $|z_{i}^{(j+1)}|<|z_{i}^{(j)}|$, until a node satisfying $z_{i}^{(k)}\in(\mathcal{B}_{A\Delta C}\cup\mathcal{B}_{AC})\setminus\left\{x\right\}$ is found. Finally, such a node can be found since if $|z_{i}^{(k)}|=1$ then $z_{i}^{(k)}$ only has two leaf descendants, one of which is in $A$ while the other is in $C$, and therefore $z_{i}^{(k)}\in\mathcal{B}_{A\Delta C}$. Overall, there exist $z_{1},z_{2}\in(\mathcal{B}_{A\Delta C}\cup\mathcal{B}_{AC})\setminus\left\{x\right\}$ which are descendants of $x_{1}$ and $x_{2}$ respectively such that $x=z_{1}\wedge z_{2}$. ∎
From the last few lemmas we deduce that the subtree induced by $\mathcal{B}_{AC}\cup\mathcal{B}_{A\Delta C}$, denoted $\mathcal{T}_{AC}^{(0)}$, is a binary tree in which $\mathcal{B}_{A\Delta C}$ are the leaves and $\mathcal{B}_{AC}$ are inner vertices (there might be other non-branching inner vertices) such that all vertices in $\mathcal{B}_{AC}$, and only them, are branching points of $\mathcal{T}_{AC}^{(0)}$. This observation allows us to construct a bijection $\tau$ as described above. Let $\langle x_{1},\dots,x_{N}\rangle$ be an ordering of $\mathcal{B}_{AC}$ satisfying $|x_{i}|\leq|x_{i+1}|$ for all $i=1,\dots,N-1$. Let’s start with $\tau(x_{1})$. By the minimality of $|x_{1}|$, it has two leaf descendants - $z_{1}^{(1)},z_{2}^{(1)}\in\mathcal{B}_{A\Delta C}$. Set $\tau(z_{1}^{(1)})=x_{1}$ and $\mathcal{T}_{AC}^{(1)}$ to be the tree induced by $\mathcal{B}_{AC}^{(1)}\cup\mathcal{B}_{A\Delta C}^{(1)}$ where we set $\mathcal{B}_{AC}^{(1)}=\mathcal{B}_{AC}\setminus\left\{x_{1}\right\}$ and $\mathcal{B}_{A\Delta C}^{(1)}=\mathcal{B}_{A\Delta C}\setminus\left\{z_{1}^{(1)}\right\}$. This means that the vertices of $\mathcal{T}_{AC}^{(1)}$ no longer match the original branching pattern, but this does not affect our construction of $\tau$. Notice that the new tree still satisfies all the previous statements with the new $\mathcal{B}_{AC}^{(1)}$ and $\mathcal{B}_{A\Delta C}^{(1)}$. We can continue this process, suppose we set $\tau(z_{1}^{(1)})=x_{1},\dots,\tau(z_{1}^{(k)})=x_{k}$, and the corresponding $\mathcal{T}_{AC}^{(k)}$, $\mathcal{B}_{AC}^{(k)}$ and $\mathcal{B}_{A\Delta C}^{(k)}$. As before, by the minimality of $|x_{k+1}|$ in $\left\{x_{k+1},\dots,x_{N}\right\}$ it has two leaf descendants in $\mathcal{T}_{AC}^{(k)}$; $z_{1}^{(k+1)},z_{2}^{(k+1)}\in\mathcal{B}_{A\Delta C}^{(k)}$. Define $\tau(z_{1}^{(k+1)})=x_{k+1}$ and $\mathcal{T}_{AC}^{(k+1)}$, $\mathcal{B}_{AC}^{(k+1)}=\mathcal{B}_{AC}^{(k)}\setminus\left\{x_{k+1}\right\}$ and $\mathcal{B}_{A\Delta C}^{(k+1)}=\mathcal{B}_{A\Delta C}^{(k)}\setminus\left\{z_{1}^{(k+1)}\right\}$. By construction, for all $i\in\left\{1,\dots,N\right\}$, $z_{1}^{(i)}$ is a descendant of $\tau(z_{1}^{(i)})$ in $\mathcal{T}_{AC}$ so,
$\forall i\in\left\{1,\dots,N\right\},\quad|\tau(z_{1}^{(i)})|>|z_{1}^{(i)}|$ (3.2)
Since $|\mathcal{B}_{A\Delta C}|=|\mathcal{B}_{AC}|+1$, after assigning $t$ to the remaining $z\in\mathcal{B}_{A\Delta C}$ we get the required bijection and (1.8) is implied. ∎
3.2. Existence of free energy
We begin by proving the existence and finiteness of the free energy.
Proof of Proposition 1.13.
The “if” part follows from the general statement of Proposition 1.7 by choosing $\gamma_{n}=h_{n}$ which is summable if (1.39) holds. For the other direction, by monotonicity of $h$ it is enough to consider $\epsilon$ of the form $\epsilon=2^{-m_{0}}$; $m_{0}\in\mathbb{N}_{0}$. Furthermore, since $h_{n}$ is assumed to be non-decreasing, we can minimize $\Phi_{n}$ by performing all branchings as close to the leaves as possible. That is, by choosing $\underline{b}_{n}=(2^{n-m_{0}-1},\dots,4,2,1,0,\dots,0)$. The next lower bound of $\Phi(A)$ follows from this reasoning:
$\Phi_{n}(A)\geq\sum_{\ell=0}^{n}h_{\ell}b_{\ell}(A)\geq\sum_{\ell=0}^{n-m_{0}}2^{n-m_{0}-\ell}h_{\ell}=2^{n-m_{0}}\sum_{\ell=0}^{n-m_{0}}2^{-\ell}h_{\ell}=a_{0}\sum_{\ell=0}^{n-m_{0}}2^{-\ell}h_{\ell}$
for every $A\subset\mathbb{L}^{(n)}$ of cardinality $a_{0}=2^{n-m_{0}}$. The case $a_{0}=0$ is trivial. Otherwise, since the total number of such $A$-s is
$\binom{2^{n}}{2^{-m_{0}}2^{n}}\sim{\rm e}^{a_{0}m_{0}\ln 2},$ (3.3)
it follows that $\omega(\epsilon_{0})=-\infty$ for any $m_{0}\in\mathbb{N}_{0}$, whenever the series in (1.39) diverge. ∎
3.3. Sufficient condition for a wetting transition
Henceforth we are going to assume that condition (1.39) holds. In particular, this condition implies that $2^{-n}h_{n}\to 0$ as $n\to\infty$. Therefore the canonical partition function $\omega$ will not be changed if we take $h=0$ instead of $h=h_{n}$ in the definition of $\Phi_{n}$ as a first ordering clustering function. While $\Phi_{n}$ is no longer monotone, we shall not make use of this property in the computations below. On the other hand, this redefinition will make these computations slightly simpler.
For what follows, if $\underline{b}=(b_{k})_{k=0}^{n}$ is branching pattern of the subtree $\mathcal{T}(A)$ induced by $A\subset\mathbb{L}_{0}^{(n)}$ in $\mathbb{T}^{(n)}$, then we shall write $A\sim\underline{b}$. Given such sub-tree we shall denote by $a_{\ell}$ the total population at age $\ell$ or older, so that for $l\in[0,n]$,
$a_{\ell}=1+\sum_{k=\ell+1}^{n}b_{k}\,,\;\ell=0,\dots,n\,.$ (3.4)
Notice that $a_{0}=b_{0}=|A|$. Given $n\geq 1$ and $a_{0}$, we say that a sequence $\underline{b}=\left(b_{0},\dots,b_{n}\right)$ is an admissible branching pattern for $a_{0}$ and write $\underline{b}\sim a_{0}$ if there exists $A\subset\mathbb{L}_{0}^{(n)}$ such that $|A|=a_{0}$ and $\underline{b}$ is the branching pattern corresponding to $A$. It is an admissible branching pattern if it is admissible for some $a_{0}$. It is not difficult to check that $\underline{b}=(b_{k})_{k=0}^{n}$ is admissible if and only if $b_{k}\in\mathbb{N}$ for all $k\in[0,n]$,
$a_{0}=1+\sum_{j=1}^{n}b_{j}\,,\qquad\text{and}\qquad b_{k}\leq a_{k}=1+\sum_{j=k+1}^{n}b_{j}\,,\,\,\forall k=1,\dots,n\,.$ (3.5)
We can therefore write the canonical partition function as
$\mathbb{W}_{n}(a_{0}):=\sum_{\underline{b}\sim a_{0}}\mathrm{ent}\left(\underline{b}\right){\rm e}^{-\Phi_{n}(\underline{b})}\,,$ (3.6)
where $\Phi_{n}(\underline{b})={\rm e}^{-\sum_{\ell=0}^{n}h_{\ell}b_{\ell}}$ and
$\mathrm{ent}\left(\underline{b}\right):=\#\{A\subset\mathbb{L}^{(n)}_{0}:\>A\sim\underline{b}\}\,.$ (3.7)
Lemma 3.5.
The entropy of a branching pattern $\underline{b}$ is given by
$\mathrm{ent}\left(\underline{b}\right)=\prod_{k=1}^{n}\binom{a_{k}}{b_{k}}2^{a_{k}-b_{k}}\,.$ (3.8)
Proof.
As previously stated, a first order branching pattern $\underline{b}$ completely determines the number of vertices in each layer of the induced sub-tree $a_{k}$. At each layer $k$, there are $\binom{a_{k}}{b_{k}}$ ways to choose the branching points of the level. Then, there are $2^{a_{k}-b_{k}}$ ways to choose the descendent of each non-branching vertex. These decisions can be performed independently of other layers, only adhering to the required $b_{k}$ branching points out of $a_{k}$, so the total number of subsets $A\subseteq\mathbb{L}_{0}$ such that $A\sim\underline{b}$ is the product of all terms as in (3.8). ∎
Proof of Proposition 1.14.
For $a_{0}\in[0,2^{n}]$ we may write
$\begin{split}\mathbb{W}_{n}(a_{0})=\sum_{\underline{b}\sim a_{0}}\mathrm{ent}\left(\underline{b}\right){\rm e}^{-\Phi_{n}(\underline{b})}&=\sum_{\underline{b}\sim a_{0}}\prod_{k=1}^{n}\binom{a_{k}}{b_{k}}2^{a_{k}-b_{k}}{\rm e}^{-\sum_{\ell=0}^{n}h_{\ell}b_{\ell}}\\ &=2^{n+2}e^{-a_{0}(2\ln 2+h_{0})}\sum_{\underline{b}\sim a_{0}}\prod_{k=1}^{n}\binom{a_{k}}{b_{k}}{\rm e}^{-\sum_{\ell=1}^{n}g_{\ell}b_{\ell}}\,,\end{split}$ (3.9)
where $g$ is as in (1.43) and we used summation by parts:
$\sum_{k=1}^{n}(a_{k}-b_{k})=(n+2)-2a_{0}+\sum_{k=1}^{n}kb_{k}\,.$ (3.10)
Using that
$\prod_{1}^{n}\binom{a_{k}}{b_{k}}\leq\prod_{1}^{n}\binom{a_{k-1}}{b_{k}}=\frac{a_{0}!}{b_{n}!\dots b_{1}!}\,,$ (3.11)
the right hand side of (3.9) is further upper bounded by
$2^{n+2}e^{-a_{0}(2\ln 2+h_{0})}\Big{(}\sum_{k=1}^{n}{\rm e}^{-g_{k}}\Big{)}^{a_{0}}=2^{n+2}\exp\Big{(}\!-a_{0}\Big{(}2\ln 2+h_{0}-\ln\sum_{k=1}^{n}{\rm e}^{-g_{k}}\Big{)}\Big{)}\,.$ (3.12)
Plugging in $a_{0}=\epsilon 2^{n}$ and using Theorem 1.19, we get
$J^{*}=-\lim_{\epsilon\to 0}\lim_{n\to\infty}\frac{\ln\mathbb{W}_{n}(\epsilon 2^{n})}{\epsilon 2^{n}}\geq 2\ln 2+h_{0}-\ln\sum_{k=1}^{n}e^{-g_{k}}=2\ln 2+h_{0}-\ln\kappa_{1}\,,$ (3.13)
with $\kappa_{1}$ as in the statement of the theorem. ∎
3.4. Necessary condition for a wetting transition.
The necessary condition will follow directly from Lemma 3.6 below. Recall that the sequence $g$ is related to $h$ as in (1.43).
Lemma 3.6.
Let $(h_{k})_{k=0}^{\infty}$ be a non-decreasing sequence and that (1.39) holds. Suppose that $g_{k}\geq 0$ for all $k\geq 0$. Then for all small enough $s>0$ such that $e^{s}-1$ is dyadic,
$-J^{*}\geq\ln\frac{1}{s}-\big{(}s+O(s^{2})\big{)}\hat{g}(s)-C\,.$ (3.14)
Proof.
If $\hat{g}(s_{0})=\infty$ for some $s_{0}>0$, then $\hat{g}(s)=\infty$ for all $s\leq s_{0}$, so that the statement holds trivially. We therefore henceforth assume that $\hat{g}(s)<\infty$ for all $s>0$. Let $(\epsilon_{j})_{j\geq 1}$ be a sequence which tends to 00 as $j\to\infty$. For each $j\geq 1$ and all $n$ large enough, let $\underline{b}^{(j,n)}=(b^{(j,n)}_{k})_{k=0}^{n}$ be an admissible branching pattern on $\mathbb{T}^{(n)}$ for $a_{0}^{(j)}=\epsilon_{j}2^{n}$. By Theorem 1.8 and since the sum in (3.9) is lower bounded by any of its terms, we have
$-J^{*}\geq\lim_{j\to\infty}\limsup_{n\to\infty}\,A_{n}(\underline{b}^{(j,n)})-C\,,$ (3.15)
where
$A_{n}(\underline{b}):=a_{0}^{-1}\sum_{k=1}^{n}\big{(}\ln{\binom{a_{k}}{b_{k}}-g_{k}b_{k}}\big{)}\,.$ (3.16)
It is therefore enough to find for any $s>0$ as in the statement of the lemma, sequences $(\epsilon_{j})_{j}$ and $(b^{(j,n)})_{j,n}$ such that the limit on the right hand side of (3.15) is at least the quantity on the right hand side of (3.14).
To this end, for an admissible branching pattern $\underline{b}=(b_{k})_{k=0}^{n}$, we set
$t_{k}=\frac{b_{k}}{a_{k}}\quad;\qquad k\in[1,n]\,,$ (3.17)
so that by (3.4)
$a_{k+1}=a_{k}(1+t_{k+1})^{-1}\quad,\qquad a_{k}=a_{0}\prod_{\ell=1}^{k}(1+t_{\ell})^{-1}\,,$ (3.18)
and then by Stirling’s approximation
$\ln{\binom{a_{k}}{b_{k}}}=\ln{\binom{a_{k}}{t_{k}a_{k}}}=a_{k}\varphi(t_{k})+O(\ln{a_{k}})=a_{0}\varphi(t_{k})\prod_{\ell=1}^{k}(1+t_{\ell})^{-1}+O(n)\,,$ (3.19)
where
$\varphi(t)=-t\ln{(t)}-(1-t)\ln{(1-t)}\,.$ (3.20)
Plugging this in (3.16), gives,
$A_{n}(\underline{b})\geq\sum_{k=1}^{n}\big{(}\varphi(t_{k})-g_{k}t_{k}\big{)}\prod_{\ell=1}^{k}\left(1+t_{\ell}\right)^{-1}+O(a_{0}^{-1}n^{2})\,.$ (3.21)
Now given a dyadic $t\in(0,1)$, for all $j\geq 1$ and $n$ large enough let $\underline{b}^{(j,n)}=(b^{(j,n)}_{k})_{k=0}^{n}$ be the sequence defined via (3.4) and (3.17) for
$t_{k}=\begin{cases}t&1\leq k\leq j\,,\\ 1&j<k\leq n\,.\end{cases}$
Then for all $n$ large enough, $\underline{b}^{(j,n)}$ is an admissible branching pattern for $a_{0}=\epsilon_{j}2^{n}$ with $\epsilon_{j}:=((1+t)/2)^{j}$. Clearly $\epsilon_{j}\to 0$ as $j\to\infty$. At the same time, by (3.21),
$\begin{split}A_{n}(\underline{b}^{(j,n)})&\geq\sum_{k=1}^{j}(1+t)^{-k}\left(\varphi(t)-tg_{k}\right)-\left(\frac{2}{1+t}\right)^{j}\sum_{k=j+1}^{n}2^{-k}g_{k}+o(1)\,,\\ &\geq\varphi(t)\sum_{k=1}^{j}(1+t)^{-k}-t\sum_{k=1}^{j}(1+t)^{-k}g_{k}-\sum_{k=j+1}^{n}(1+t)^{-k}g_{k}+o(1)\,,\end{split}$ (3.22)
with $o(1)\to 0$ as $n\to\infty$ for all $j$. Finiteness of the Laplace transform of $g$ for all $s>0$ shows that the last sum tends to 00 when $n\to\infty$ followed by $j\to\infty$. Taking these limits, we thus get, in light of (3.15), whenever $t$ is small enough,
$-J^{*}\geq\frac{\varphi(t)}{t}-t\hat{g}\big{(}\ln(1+t)\big{)}-C\geq\ln\frac{1}{t}-t\hat{g}\big{(}\ln(1+t)\big{)}-C^{\prime}\,.$ (3.23)
where we have used that $\varphi(t)/t=\ln{\frac{1}{t}}+O(1)$ near 00. Finally, substituting $t=e^{s}-1=s+O(s^{2})$ as $s\to 0$ and noting that $-\ln(s+O(s^{2}))=-\ln s+O(s)$ in this regime, this gives the desired statement. ∎
We immediately get
Proof of Proposition 1.15.
Since the free energy only decreases if we replace $h_{k}$ by $h_{k}\vee(\ln 2)k$, which is also non-decreasing and satisfies (1.39), we may prove the statement under the additional assumption that $g_{k}\geq 0$. Then Lemma 3.6 is in force, and (1.42) gives $|s\hat{g}(s)|=s\widehat{g^{+}}(s)=O(\ln(1/s))$, so that $O(s^{2})\hat{g}(s)\to 0$ as $s\to 0$. It then follows from Lemma 3.6 that $-J^{*}=\infty$. ∎
Proof of Corollary 1.16.
As in in the proof of Proposition 1.15, we shall assume w.l.o.g. that $g$ is non-negative. It follows from [9, Corollary 1c], and that the Laplace transform of the (possibly negative) sequence $(u_{n}-u_{n-1})_{n\geq 1}$ is $(1-e^{-s})\hat{u}(s)+u_{0}\leq 2s\hat{u}(s)+u_{0}$, that
$\liminf_{s\downarrow 0}2s\hat{u}(s)\geq\liminf_{k\to\infty}u_{k}-u_{0}\,,$ (3.24)
for any sequence $u=(u_{n})_{n\geq 0}$ for which $\hat{u}(s)<\infty$ for all $s>0$. Now suppose that (1.45) holds and take $u=(u_{k})_{k\geq 0}$ to be sequence defined via $u_{k}:=\ln(k\vee 1)-g_{k}$. Then by positivity of $g$ and (1.45) we have $\hat{u}(s)<\infty$ for all $s>0$. Noting that the Laplace transform of $\ln(k\vee 1)$ is $s^{-1}\ln s^{-1}+O(s^{-1})$ as $s\to 0$, we thus get from (3.24)
$\liminf_{s\downarrow 0}\big{(}\ln\frac{1}{s}-s\hat{g}(s)\big{)}\geq\frac{1}{2}\liminf_{k\to\infty}(\ln k-g_{k})-C=\infty\,,$ (3.25)
so that there is no wetting transition by Proposition 1.15. ∎
4. Clustering via second order branching patterns
4.1. Monotoncity
Proof of Proposition 1.18.
The proof of (1.7) will be done by induction on the cardinality $N+1$ of $A\subseteq\mathbb{L}_{0}$. (1.7) is trivial for $N=1$. Assume that (1.7) holds for cardinality $N+1$ and let $A\prec B$ be two subsets of $\mathbb{L}_{0}$ of cardinality $N+2$.
We shall label leaves in $A$ as $21,\dots,N+$ 2 and leaves in $B$ as $\sigma_{1},\dots,\sigma_{N+2}$, where $\sigma$ is the correspondence which realizes $A\prec B$, that is $|i\wedge j|\leq|\sigma_{i}\wedge\sigma_{j}|$ for any $i\neq j$. Define
$m_{i}=\min_{j\neq i}|i\wedge j|,\ i^{*}=\min\left\{j\leavevmode\nobreak\ :\leavevmode\nobreak\ |i\wedge j|=m_{i}\right\}\ \text{ and }\ r_{i}=\min_{j\neq i,i^{*}}|i\wedge j|\,.$ (4.1)
Without loss of generality we may assume that $|1\wedge 2|=m_{1}=\min_{i}m_{i}$. Note that this necessarily implies that $r_{1}>m_{1}$ (otherwise $m_{2}<m_{1}$). Consider now $\hat{A}=A\setminus\left\{1\right\}$ and $\hat{B}=B\setminus\{\sigma_{1}\}$. Clearly, $\hat{A}\prec\hat{B}$, and hence by the induction hypothesis $\Phi_{n}(\hat{A})\leq\Phi_{n}(\hat{B})$. By construction,
$\Phi_{n}(A)=\Phi_{n}(\hat{A})+h_{r_{1},m_{1}}.$ (4.2)
In order to compare (4.2) with the corresponding contribution to $\Phi_{n}(B)$ we shall consider two cases.
CASE 1. $r_{\sigma_{1}}>m_{\sigma_{1}}$. Then there exists unique $j$ such that $m_{\sigma_{1}}=|\sigma_{1}\wedge\sigma_{j}|$. It follows that $m_{\sigma_{1}}\geq m_{1}$ and $r_{\sigma_{1}}\geq r_{1}$. We are adding a branching point in generation $m_{\sigma_{1}}$ with branching ancestor in generation $r_{\sigma_{1}}$. Hence,
$\Phi_{n}(B)=\Phi_{n}(\hat{B})+h_{r_{\sigma_{1}},m_{\sigma_{1}}}\geq\Phi_{n}(\hat{A})+h_{r_{1},m_{1}}=\Phi_{n}(A).$ (4.3)
CASE 2. $r_{\sigma_{1}}=m_{\sigma_{1}}$. In this case we add a branching point in generation $r_{\sigma_{1}}\geq r_{1}$. So its branching ancestor belongs to generation $r$ with $r>r_{1}$. Since $r_{\sigma_{1}}=m_{\sigma_{1}}$ it also has a branching descendant in generation $m$ with $m\geq m_{1}$. Which means that adding $\sigma_{1}$ to $\hat{B}$ removes a branching point of type $(r,m)$ and instead adds two branching points - one of type $(r_{\sigma_{1}},m)$ and one of type $(r,r_{\sigma_{1}})$. Therefore,
$\Phi_{n}(B)$ $=\Phi_{n}(\hat{B})+\left(h_{r,r_{\sigma_{1}}}-h_{r,m}\right)+h_{r_{\sigma_{1}},m}\,\geq\,\Phi_{n}(\hat{B})+h_{r_{\sigma_{1}},m}$ (4.4)
$\geq\Phi_{n}(\hat{A})+h_{r_{1},m_{1}}\,=\,\Phi_{n}(A)$ (4.5)
Let us turn to verifying (1.8). Let $A,B$ be two disjoint subsets of $\mathbb{L}_{0}$.
Consider first $\mathcal{B}(A)\cap\mathcal{B}(B)=\emptyset$. Then there exists unique $z$ such that
$\left\{{z}\right\}=\operatorname*{arg\,min}\left\{\left|u\wedge v\right|\leavevmode\nobreak\ :\leavevmode\nobreak\ u\in A,v\in B\right\}.$
Therefore, $\mathcal{B}(A\cup B)=\mathcal{B}(A)\cup\mathcal{B}(B)\cup\left\{z\right\}$. For each $u\in\mathcal{B}(A\cup B)$, let $\mathfrak{a}_{A\cup B}(u)$ be its branching ancestor.
Remark 4.1.
Recall that by our convention there is always a branching ancestor in some generation $\ell\leq n+1$. In the sequel we shall, with a certain abuse of notation, use $\left|\mathfrak{a}_{A}(u)\right|$ for the generation of $\mathfrak{a}_{A}(u)$. In this way it might happen that $\left|\mathfrak{a}_{A}(u)\right|=n+1$.
Now, we can rewrite (1.50) as follows
$\Phi_{n}(A\cup B)=\sum_{u\in(A\cup B)\setminus\left\{z\right\}}h_{\left|\mathfrak{a}_{A\cup B}(u)\right|,\left|u\right|}+h_{\left|\mathfrak{a}_{A\cup B}(z)\right|,\left|z\right|}+h.$ (4.6)
For any $u\in\mathcal{B}(A)$, $\left|\mathfrak{a}_{A}(u)\right|\geq\left|\mathfrak{a}_{A\cup B}(u)\right|$. The same regarding $u\in\mathcal{B}(B)$. Hence, the first term on the right hand side of (4.6) is less or equal to
$\sum_{u\in\mathcal{B}(A)}h_{\left|\mathfrak{a}_{A}(u)\right|,\left|u\right|}+\sum_{u\in\mathcal{B}(B)}h_{\left|\mathfrak{a}_{B}(u)\right|,\left|u\right|}.$
Since, by construction, $h_{\left|\mathfrak{a}_{A\cup B}(z)\right|,\left|z\right|}\leq h$, the assertion $\Phi(A\cup B)\leq\Phi(A)+\Phi(B)$ follows.
If $\left|\mathcal{B}(A)\cap\mathcal{B}(B)\right|=N>0$, then the following happens:
  • As before for any $u\in\mathcal{B}(A)$, it holds that $h_{\left|\mathfrak{a}_{A}(u)\right|,\left|u\right|}\geq h_{\left|\mathfrak{a}_{A\cup B}(u)\right|,\left|u\right|}$. The same regarding $u\in\mathcal{B}(B)$.
  • However, each one of $N$ branching points $w\in\mathcal{B}_{AB}:=\mathcal{B}(A)\cap\mathcal{B}(B)$ is counted twice in $\Phi_{n}(A)+\Phi_{n}(B)$, whereas it is counted once in $\Phi_{n}(A\cup B)$.
  • On the other hand there are exactly $N+1$ new branching points $z\in\mathcal{B}_{A\Delta B}:=\mathcal{B}(A\cup B)\setminus\left(\mathcal{B}(A)\cup\mathcal{B}(B)\right)$.
Therefore we just need to find a matching $\tau$ between any $N$ out of the $(N+1)$ points of $\mathcal{B}_{A\Delta B}$ and $N$ points of $\mathcal{B}_{AB}$ such that,
$h_{\left|\mathfrak{a}_{A}(\tau(z))\right|\wedge\left|\mathfrak{a}_{B}\tau(z))\right|,\left|\tau(z)\right|}\geq h_{\left|\mathfrak{a}_{A\cup B}(z)\right|,\left|z\right|}$ (4.7)
for any one of these $N$ chosen points $z\in\mathcal{B}_{A\Delta B}$. Indeed, given such a $\tau$ the remaining $(N+1)^{\text{\text@underline{st}}}$ point $w$ of $\mathcal{B}_{A\Delta B}$ will always satisfy $h_{\left|\mathfrak{a}_{A\cup B}(w)\right|,\left|w\right|}\leq h$, and hence (4.7) indeed implies (1.8) in its full generality. The mapping $\tau$ defined in the previous subsection applies here as well. For all $z\in\mathcal{B}_{A\Delta C},z\neq w$
$|\mathfrak{a}_{A\cup B}(z)|\leq|\tau(z)|<|\mathfrak{a}_{A\cup B}(\tau(z))|\leq|\mathfrak{a}_{A}(\tau(z))|\wedge|\mathfrak{a}_{B}(\tau(z))|$
This completes the proof. ∎
4.2. Existence of free energy
Proof of proposition 1.19.
Let $n\in\mathbb{N}$ and $A\subseteq\mathbb{L}_{0,r/l}^{(n)}$. Denote by $\ell^{*}$ the maximal height of a branching point in the tree induced by $A$. Then for $A\neq\emptyset$, by monotonicity of $\{h_{k,\ell}\}$,
$\Phi_{n}(A)-\Phi_{n-1}(A)$ $=h_{n+1,n}+h_{n+1,\ell^{*}}-h_{n,n-1}-h_{n,\ell^{*}}$
$\leq 2h_{n+1,n}\,.$
Set $\gamma_{n}=2h_{n+1,n}$ then $\sum_{n}\gamma_{n}2^{-n}<\infty$ and assumption (1.14) holds. By Proposition 1.7 we get the desired conclusion. ∎
4.3. Sufficient condition for wetting
As in the case of first order branching, for the sake of simplifying the proofs to follow, we shall assume henceforth that $h=0$ instead of $h_{n+1,n}$ in the definition of $\Phi_{n}$ as a second order branching clustering function. This will again not change the canonical free energy as $2^{-n}h_{n+1,n}\to 0$ as $n\to\infty$ thanks to the assumption (1.52) which will be assumed throughout.
Also, as before, if $\underline{b}$ is the second order branching pattern for the subtree $\mathcal{T}(A)$ induced by $A\subset\mathbb{L}_{0}^{(n)}$ in $\mathbb{T}^{(n)}$, we write $A\sim\underline{b}$. For $\ell\in[0,n]$ we also let $a_{\ell}$ be the total population size at generation $\ell$ or higher, so that
$a_{\ell}:=1+\sum_{j=\ell}^{n}b_{j}\;,\quad\ell=1,\dots,n\;,\quad a_{0}=b_{0}=|A|\,,$ (4.8)
where we recall that $(b_{k})_{k=0}^{n}$ is the first order branching pattern, which is related to the second order branching pattern via,
$2b_{k}=\sum_{\ell=0}^{k-1}b_{k,\ell}\;,\quad b_{\ell}=\sum_{k=\ell+1}^{n+1}b_{k,\ell}\,,$ (4.9)
for all $\ell,k\in[0,n]$.
Given $n\geq 1$, $a_{0}\in[0,2^{n}]$, we shall say that a triangular array $\underline{b}=(b_{k,\ell})_{0\leq\ell<k\leq n+1}$ is an admissible second order branching pattern for $a_{0}$ and write $\underline{b}\sim a_{0}$ if there exists a subset $A\subseteq\mathbb{L}^{(n)}_{0}$ with $|A|=a_{0}$, such that $A\sim\underline{b}$. It is an admissible second order branching pattern if it is admissible for some $a_{0}$. The number of subsets of leaves $A\subset\mathbb{L}_{0}^{(n)}$ with a given (admissible) second order branching pattern $\underline{b}$ is given by
$\mathrm{ent}\left(\underline{b}\right)=\#\{A\subseteq\mathbb{L}^{(n)}_{0}\,:\,A\sim\underline{b}\}\,.$ (4.10)
As in the first order branching case, we can write the canonical partition function as
$\mathbb{W}_{n}(a_{0}):=\sum_{\underline{b}\sim a_{0}}\mathrm{ent}\left(\underline{b}\right){\rm e}^{-\Phi_{n}(\underline{b})}\,,$ (4.11)
where
$\Phi_{n}(\underline{b})=\sum_{k=1}^{n+2}\sum_{\ell=0}^{k-1}h_{k,\ell}b_{k,\ell}\,,$ (4.12)
with the convention that $b_{n+2,\ell}={1\hskip-3.0pt\rm{I}}_{\left\{\ell=n+1\right\}}$.
Lemma 4.2.
For $n\geq 1$, if $\underline{b}=(b_{k,\ell})_{0\leq\ell<k\leq n+1}$ is an admissible second order branching pattern, then
$\mathrm{ent}\left(\underline{b}\right)=\prod_{j=1}^{n}\binom{2b_{j}}{b_{j,j-1},\ldots,b_{j,0}}2^{a_{j}-b_{j}},$ (4.13)
above $\binom{2b_{j}}{b_{j,j-1},\ldots,b_{j,0}}:=b_{j}!\Big{/}b_{j,j-1}!\dots b_{j,1}!b_{j,0}!$ is the multinomial coefficient.
Proof.
Notice that two sets $A$ and $A^{\prime}$ that are compatible with $\underline{b}$ may be different in a graph sense, nevertheless they will always have the same population size in every generation. Each non branching point has exactly one offspring that can be embedded in $\mathbb{T}_{n}$ in two distinct ways (“right” or “left”). Therefore, after fixing the branching points in generation $j$ there are $2^{a_{j}-b_{j}}$ different ways to obtain the next vertex in generation $j-1$, giving the $2^{a_{j}-b_{j}}$ term in (4.13). Finally the first term in the product is the number of possible ways one may arrange the branching points $b_{j+i,j}$ in generation $j+i$ whose last branching ancestor is in generation $j$. ∎
The next lemma shows that the free energy does not change if one replaces the sum in (4.11) by the maximum.
Lemma 4.3.
For all dyadic $\epsilon>0$,
$\omega(\epsilon)=\lim_{n\to\infty}\frac{1}{2^{n}}\ln\max_{\underline{b}\sim\epsilon 2^{n}}\big{(}\mathrm{ent}\left(\underline{b}\right){\rm e}^{-\Phi_{n}(\underline{b})}\big{)}\,.$ (4.14)
Proof.
Clearly,
$\max_{\underline{b}\sim\epsilon 2^{n}}\big{(}\mathrm{ent}\left(\underline{b}\right){\rm e}^{-\Phi_{n}(\underline{b})}\big{)}\leq\,\mathbb{W}_{0}(\epsilon 2^{n})\,\leq\#\{\underline{b}:\>\underline{b}\sim\epsilon 2^{n}\}\max_{\underline{b}\sim\epsilon 2^{n}}\big{(}\mathrm{ent}\left(\underline{b}\right){\rm e}^{-\Phi_{n}(\underline{b})}\big{)}$ (4.15)
Since $\sum_{0\leq\ell<k\leq n+1}b_{k,\ell}=2a_{0}-1=2\epsilon 2^{n}-1$, the total number of admissible arrays $\underline{b}\sim\epsilon 2^{n}$ is smaller than $(2\epsilon 2^{n}-1)^{n^{2}}\leq{\rm e}^{Cn^{3}}$. Therefore, the limit as $n\to\infty$ of $2^{-n}$ times the logarithm of both bounds in (4.15) tend to the right hand side of (4.14). ∎
For all $k\in\mathbb{N}$ and $10\leq\ell\leq k-$ 1, set
$p_{k,\ell}=h_{k,\ell}-(\ln 2)\ell\,.$ (4.16)
Proof of Proposition 1.20.
Let $\epsilon>0$, $n\geq 1$ and take $\underline{b}\sim\epsilon 2^{n}$ to be the maximizer in (4.14). As in (3.10) we can wrtie
$\begin{split}\sum\limits_{j=1}^{n}(a_{j}-b_{j})&=(n+2)-2a_{0}+\sum\limits_{\ell=0}^{n}\ell b_{\ell}\\ &=(n+2)-2a_{0}+\sum\limits_{\ell=0}^{n}\sum\limits_{k=\ell+1}^{n+1}\ell b_{k,\ell}=(n+2)-2a_{0}+\sum\limits_{k=1}^{n+1}\sum\limits_{\ell=0}^{k-1}\ell b_{k,\ell}\end{split}$ (4.17)
Thanks to Lemma 4.2 and Lemma 4.3 we thus get
$\begin{split}\omega(\epsilon)&=-\epsilon(2\log 2)+\lim_{n\to\infty}2^{-n}\log\left(\left[\prod\limits_{k=1}^{n}\binom{2b_{k}}{b_{k,k-1},\ldots,b_{k,0}}\right]\times\prod\limits_{k=1}^{n+2}\prod\limits_{\ell=0}^{k-1}e^{-p_{k,\ell}b_{k,\ell}}\right)\\ &=-\epsilon(2\log 2)+\lim_{n\to\infty}2^{-n}\sum_{k=1}^{n}\log\left[\binom{2b_{k}}{b_{k,k-1},\ldots,b_{k,0}}\prod_{\ell=0}^{k-1}e^{-p_{k,\ell}b_{k,\ell}}\right]\,.\end{split}$ (4.18)
where the last equality follows since
$\begin{split}\sum_{k=n+1}^{n+2}&\sum_{\ell=0}^{k-1}p_{k,\ell}b_{k,\ell}\leq\max_{k=n+1,n+1}\max_{\ell=0,\dots,k-1}\left|p_{k,\ell}\right|\sum_{k=n+1}^{n+2}\sum_{\ell=0}^{k-1}b_{k,\ell}\\ &=\frac{1}{2}\left(b_{n+1}+b_{n+2}\right)\max_{k=n+1,n+2}\max_{\ell=0,\dots,k-1}\left|p_{k,\ell}\right|\leq\left(h_{n+2,n+1}+\left(n+1\right)\ln{2}\right)\,.\end{split}$ (4.19)
By Stirling’s approximation
$\log\binom{2b_{k}}{b_{k,k-1}\ \ldots\ b_{k,0}}=2b_{k}\log 2b_{k}-\sum_{l=0}^{k-1}\Big{(}b_{k,l}\log b_{k,l}\Big{)}+O(n^{2})\,,$ (4.20)
so the last sum in (4.18) is equal to $O(n^{3})$ plus
$\begin{split}-\sum_{k=1}^{n}\sum_{\ell=0}^{k-1}b_{k,\ell}\log\Big{(}\frac{b_{k,\ell}}{2b_{k}e^{-p_{k,\ell}}}\Big{)}&=-\sum_{k=1}^{n}2b_{k}\left[\sum_{\ell=0}^{k-1}e^{-p_{k,\ell}}\Big{(}\frac{b_{k,\ell}}{2b_{k}e^{-p_{k,\ell}}}\Big{)}\log\Big{(}\frac{b_{k,\ell}}{2b_{k}e^{-p_{k,\ell}}}\Big{)}\right]\\ &\leq e^{-1}\sum_{k=1}^{n}2b_{k}\sum_{\ell=0}^{k-1}e^{-p_{k,\ell}}\leq 2e^{-1}\kappa_{2}\epsilon 2^{n}\,.\end{split}$ (4.21)
Above, the first inequality follows since $y\log y\geq-e^{-1}$ for all $y>0$, and $\kappa_{2}$ is as in (1.53). Plugging this in (4.18), taking the limit as $\epsilon\to 0$ and using Theorem 1.8 we complete the proof. ∎
4.4. Necessary condition for wetting
Recall the definition of $g_{\ell,d}$ from (1.58).
Lemma 4.4.
Let $(h_{k,j})_{0\leq j<k}$ be non-decreasing triangular array and suppose that (1.52) holds. Suppose also that $g_{\ell,d}\geq 0$ for all $\ell,d\geq 0$. Then there exists arbitrarily small $s$ such that $e^{s}-1$ is dyadic and
[…]
Proof.
We proceed as in the case of first order branching. If $\widehat{g}(s,2s)=\infty$ for some $s>0$, then the statement holds trivially, so we shall assume otherwise. Let $(\epsilon_{j})_{j\geq 1}$ be a sequence which tends to 00 as $j\to\infty$. For each $j\geq 1$ and all $n$ large enough, let $\underline{b}^{(j,n)}=(b^{(j,n)}_{k,\ell})_{0\leq\ell<k\leq n}^{n}$ be an admissible second order branching pattern on $\mathbb{T}^{(n)}$ for $a_{0}^{(j,n)}=\epsilon_{j}2^{n}$. It follows from Theorem 1.8 and (4.18) that
$-J^{*}\geq\lim_{j\to\infty}\limsup_{n\to\infty}\,B_{n}(\underline{b}^{(j,n)})-C\,,$ (4.23)
where
$B_{n}(b)=a_{0}^{-1}\sum_{k=1}^{n}\ln\left[\binom{2b_{k}}{b_{k,k-1},\ldots,b_{k,0}}\prod_{d=1}^{k}e^{-g_{k-d,d}b_{k,k-d}}\right]$ (4.24)
Now, for $t>0$ dyadic, $j\in\mathbb{N}$ and all $n\geq j$ set,
$t_{k}:=\begin{cases}1&j<k<n\,,\\ t&0\leq k\leq j\,,\end{cases}$ (4.25)
let $(a_{k})_{k=0}^{n}$, $(b_{k})_{k=0}^{n}$ be defined via (4.8),
[…]
and set
$b_{k,k-d}=\begin{cases}2b_{k}\cdot t_{k-1}\cdot(1-t_{k-1})^{d-1}&1\leq d<k\leq n\,,\\ 2b_{k}\cdot(1-t_{k-1})^{k-1}&d=k\in[1,n]\,.\end{cases}$ (4.27)
We claim that for all such $t$ and $j$ whenever $n$ is large enough the triangular array $b=(b_{k,\ell})_{0\leq\ell<k\leq n+1}$ is an admissible second order branching pattern for $a_{0}=\epsilon_{j}2^{n}$ with $\epsilon_{j}=((1+t)/2)^{j}$. To see this, we need to verify that the $b_{k,\ell}$-s are all integer and that (4.9) holds. Indeed,
$b_{k}=\begin{cases}2^{n-k}&j+1\leq k\leq n\,,\\ 2^{n-j}(1+t)^{j-k}t&1\leq k\leq j\,,\end{cases}$ (4.28)
and
$b_{k,k-d}=\begin{cases}2b_{k}1_{\{d=1\}}&k\geq j+1\,,\\ 2b_{k}t(1-t)^{d-1}&1\leq d<k\leq j\,,\\ 2b_{k}(1-t)^{k-1}&d=k\leq j\,.\end{cases}$ (4.29)
The right hand side above can be made integer for all dyadic $t>0$ and $j\in\mathbb{N}$ by taking $n$ large enough. As for (4.9), the first equality in (4.9) follows easily by summing over $d=1,\dots,k$ in (4.29). The second trivially holds for $\ell\geq j+1$. For $1\leq\ell\leq j$, we have
$\begin{split}\sum_{k=\ell+1}^{n+1}b_{k,\ell}&=\sum_{d=1}^{j-\ell}2b_{\ell+d}t(1-t)^{d-1}=\sum_{d=1}^{j-\ell}2\cdot 2^{n-j}t(1+t)^{j-\ell-d}\cdot t(1-t)^{d-1}\\ &=2t^{2}\cdot 2^{n-j}(1+t)^{j-\ell-1}\cdot\frac{1-\left(\frac{1-t}{1+t}\right)^{j-\ell}}{1-\frac{1-t}{1+t}}=2^{n-j}\cdot t(1+t)^{j-\ell}=b_{\ell}\,,\end{split}$ (4.30)
Turning to $B_{n}(\underline{b})$, for the sum in (4.24) when $k\in[j+1,n]$ we have,
$a_{0}^{-1}\sum_{k=j+1}^{n}2b_{k}g_{k-1,1}=\left(\frac{2}{1+t}\right)^{j}2^{-n}\sum_{k=j+1}^{n}2\cdot 2^{n-k}g_{k-1,1}\leq\sum_{\ell=j}^{n-1}(1+t)^{-\ell}g_{\ell,1}\,,$ (4.31)
since $\hat{g}(s_{1},s_{2})$ is finite for all $(s_{1},s_{2})\in(0,\infty)^{2}$, the last sum tends 00 as $n\to\infty$ followed by $j\to\infty$.
At the same time, as in (4.21), the sum in (4.24) taken over $k\leq j$, is equal to $O(n^{3})$ plus
$\begin{split}-2\sum_{k=1}^{j}b_{k}\sum_{d=1}^{k}&\frac{b_{k,k-d}}{2b_{k}}\ln{\frac{b_{k,k-d}}{2b_{k}e^{-g_{k-d,d}}}}\\ &=-2\sum_{k=1}^{j}b_{k}\sum_{d=1}^{k}t(1-t)^{d-1}\left((d-1)\ln{(1-t)}+\ln{t}+g_{k-d,d}\right)\\ &\geq-2\sum_{k=1}^{j}b_{k}\sum_{d=1}^{k}t(1-t)^{d-1}\left(\ln{t}+g_{k-d,d}\right)\\ &=-2a_{0}t^{2}\sum_{\ell=0}^{j-1}\sum_{d=1}^{j-\ell}(1+t)^{-\ell-d}(1-t)^{d-1}\left(\ln{t}+g_{\ell,d}\right)\,,\end{split}$ (4.32)
where for the last equality we plugged in $b_{k}=ta_{k}=a_{0}t(1+t)^{-k}$ and substituted $\ell=k-d$. Thanks to the assumed positivity of $g$, taking $j\to\infty$ amounts to turning the last double sum into a double series with both indices going to $\infty$. This gives
$\frac{1}{1-t}\sum_{\ell=0}^{\infty}\sum_{d=1}^{\infty}(1+t)^{-\ell}\left(\frac{1+t}{1-t}\right)^{-d}(\ln t+g_{\ell,d})\leq\frac{1}{2t^{2}}\ln t+\sum_{\ell,d=0}^{\infty}(1+t)^{-\ell}\left(1+t\right)^{-2d}g_{\ell,d}\,,$ (4.33)
where we have used the explicit formula for the geometric series and the positivity of $g$. Collecting all estimates, the double limit in (4.23) is at least
[…]
Finally, substituting $t=e^{s}-1=s+O(s^{2})$ and noting that $\ln t^{-1}=\ln s^{-1}+O(s)$, we get the desired statement. ∎
We can now give
Proof of Proposition 1.21.
By replacing $g_{\ell,d}$ by $g^{+}_{\ell,d}=g_{\ell,d}\vee 0\geq g_{\ell,d}$, which only decreases the free energy, we may assume w.l.o.g. that $g_{\ell,d}\geq 0$ for all $\ell$ and $d$. Then (1.55) implies that $\hat{g}(s,2s)=O(s^{-2}\log(1/s))$ so that $\hat{g}(s,2s)O(s^{3})=o(1)$ as $s\to 0$. In view of Lemma 4.4 we thus get $J^{*}=-\infty$ and therefore there is no wetting transition. ∎
Proof of Corollary 1.22.
Define $g^{(1)}_{l}:=h^{(1)}_{l}-(\ln 2)l$, so that
$g_{\ell,d}=g^{\left(1\right)}_{\ell}+h^{\left(2\right)}_{d}\,.$ (4.35)
By assumption $|g^{(1)}_{k}|,|h_{k}^{(2)}|=O(k)$ as $k\to\infty$, so that $|\hat{g}^{(1)}(s)|,|\hat{h}^{(2)}(s)|=O(s^{-2})$ as $s\downarrow 0$. The above additive structure then gives
$\begin{split}\hat{g}\left(s,2s\right)&=\sum_{\ell=0}^{\infty}\sum_{d=0}^{\infty}e^{-s\ell-2sd}\left(g^{\left(1\right)}_{\ell}+h^{\left(2\right)}_{d}\right)\\ &=\sum_{\ell=0}^{\infty}e^{-s\ell}g^{\left(1\right)}_{\ell}\sum_{d=0}^{\infty}\left(e^{-2s}\right)^{d}+\sum_{d=0}^{\infty}e^{-2sd}h^{\left(2\right)}_{d}\sum_{\ell=0}^{\infty}\left(e^{-s}\right)^{\ell}\\ &=\left(\frac{1}{2s}\hat{g}^{\left(1\right)}\left(s\right)+\frac{1}{s}\hat{h}^{\left(2\right)}\left(2s\right)\right)\left(1+O\left(s\right)\right)\\ &=\frac{1}{2s}\hat{g}^{\left(1\right)}\left(s\right)+\frac{1}{s}\hat{h}^{\left(2\right)}\left(2s\right)+O(s^{-2})\,.\end{split}$ (4.36)
At the same time,
$\hat{h}_{\lfloor\cdot/2\rfloor}^{\left(2\right)}(s)$ $=\sum_{d=0}^{\infty}e^{-ds}h_{\lfloor\frac{d}{2}\rfloor}^{\left(2\right)}=(2+O(s))\sum_{d=0}^{\infty}e^{-2ds}h_{d}^{\left(2\right)}=2\hat{h}^{\left(2\right)}(2s)+O(s^{-1})\,.$ (4.37)
Now, as in the proof of Corollary 1.16, condition (1.60) implies that
$\lim_{s\downarrow 0}\ln{\frac{1}{s}}-s\hat{g}^{\left(1\right)}\left(s\right)-s\hat{h}_{\lfloor\cdot/2\rfloor}^{\left(2\right)}\left(s\right)=\infty\,,$ (4.38)
where we have used the linearity of the Laplace Transform. Combining (4.36), (4.37) and (4.38), this gives
$\lim_{s\downarrow 0}\left(\ln{\frac{1}{s}}-2s^{2}\hat{g}\left(s,2s\right)\right)=\infty$
so there is no wetting transition. ∎
5. Clustering via capacity
5.1. Properness of definition
Proof of Proposition 1.24.
The proof is based on the following argument by Lyons [6, 7]: Let $f$ be the minimizer ($\mathbb{T}^{(n)}$ is a finite graph - no ambiguity) in (1.61). Then
$\theta_{A}(e):=\frac{1}{{\rm CAP}_{\mathcal{C}}(A)}\mathsf{C}_{e}\nabla_{e}f$ (5.1)
is a unit flow from $\mathsf{0}$ to $\mathbb{L}_{0}$. Define resistances $\mathsf{R}_{e}=\mathsf{C}_{e}^{-1}$. Given a flow $\theta$ define its energy $\mathcal{E}(\theta)=\sum_{e}\mathsf{R}_{e}\theta(e)^{2}$. Then, $\mathcal{E}(\theta_{A})={\rm CAP}_{\mathcal{C}}(A)^{-1}$, and $\theta_{A}$ has the minimal energy within the family $\mathcal{F}_{A}$ of unit flows from $\mathsf{0}$ to $A$.
Next, any unit flow $\theta\in\mathcal{F}_{A}$ could be viewed as a probability distribution $\mu$ on $A\subseteq\mathbb{L}_{0}^{(n)}$; $\mu(u)=\theta(e(u))$. Kirchoff’s law implies that for any $u\not\in\left\{\mathsf{0}\right\}\cup\mathbb{L}_{0}$, $\theta(e(u))=\sum_{e^{\prime}}\theta(e^{\prime}),$ where the summation is with respect all outgoing edges from $u$ in the direction of $\mathbb{L}_{0}$. Consequently, if $\mathcal{D}(u)\subseteq\mathbb{L}_{0}$ is the set of all the descendants of $u$ in $\mathbb{L}_{0}$, then $\theta(e(u))=\sum_{v\in\mathcal{D}(u)}\mu(v).$
Recall that $R_{e}=R_{\left|e\right|}$, and set
$\alpha_{\ell}^{(n)}={1\hskip-3.0pt\rm{I}}_{\ell<n}\sum_{j=\ell}^{n}\mathsf{R}_{j}.$ (5.2)
Then, summing by parts in
$\mathcal{E}(\theta)=\sum_{\left|u\right|<n}R_{\left|u\right|}\left(\sum_{v\in\mathcal{D}(u)}\mu(v)\right)^{2}$
one recovers the following formula:
$\mathcal{E}(\theta)=\sum_{u,v}\mu(u)\mu(v)\alpha_{\left|u\wedge v\right|}^{(n)}\ \text{ and }\ {\rm CAP}_{\mathcal{C}}(A)=\max_{\mu(A)=1}\left\{\sum_{u,v}\mu(u)\mu(v)\alpha_{\left|u\wedge v\right|}^{(n)}\right\}^{-1}$ (5.3)
At this point, since $\alpha^{(n)}$ in (5.2) is non-increasing, we recover (1.7), that is $A\prec B$ implies that ${\rm CAP}_{\mathcal{C}}(A)\leq{\rm CAP}_{\mathcal{C}}(B)$. ∎
5.2. No wetting transition
Proof.
Since we know that ${\rm CAP}_{\mathcal{C}}$ is a monotone clustering function,
[…]
It follows that $\mathbb{Z}^{\Phi,-J}_{n}$ in (1.10) with $\Phi(A)={\rm CAP}_{\mathcal{C}}(A)$, satisfies
$\mathbb{Z}_{n}^{\Phi,-J}\geq\sum_{N=1}^{2^{n}}\binom{2^{n}}{N}\,{\rm e}^{-N(J+\mathsf{C}_{0})}.$
Since,
$\lim_{m_{0}\to\infty}\lim_{n\to\infty}\frac{1}{2^{n-m_{0}}}\ln\binom{2^{n}}{2^{n-m_{0}}}=\infty,$
the result follows from Corollary 1.9. ∎
Acknowledgments
The research was supported by ISF grants no. 2870/21 and 1723/14, and also by the BSF award 2018330.